### v5

# Distributional Finite Element curl div Complexes and Application to Quad Curl Problems

Abstract.
The paper addresses the challenge of constructing finite element ${\rm curl\,}\operatorname{div}$ complexes in three dimensions. Tangential-normal continuity is introduced in order to develop distributional finite element ${\rm curl\,}\operatorname{div}$ complexes. The spaces constructed are applied to discretize the quad curl problem, demonstrating optimal order of convergence. Furthermore, a hybridization technique is proposed, demonstrating its equivalence to nonconforming finite elements and weak Galerkin methods.
Key words and phrases:
distributional finite element ${\rm curl\,}\operatorname{div}$ complex, quad curl problem, error analysis, hybridization
2020 Mathematics Subject Classification:
58J10; 65N12; 65N22; 65N30;
The first author was partially supported by NSF DMS-2012465, DMS-2309777, and DMS-2309785. The second author was supported by the National Natural Science Foundation of China Project 12171300, and the Natural Science Foundation of Shanghai 21ZR1480500.
1. Introduction
In this work, we will construct distributional finite element curldiv complexes in three dimensions, and apply it to solve the fourth-order curl problem $-{\rm curl\,}\Delta{\rm curl\,}\boldsymbol{u}=\boldsymbol{f},\operatorname{div}\boldsymbol{u}=0$ in a domain $\Omega\subset\mathbb{R}^{3}$ with boundary conditions $\boldsymbol{u}\times\boldsymbol{n}=\operatorname{curl}\boldsymbol{u}=0$ on $\partial\Omega$. Such problem arises from multiphysics simulation such as modeling a magnetized plasma in magnetohydrodynamics [31, 9].
We first give a brief literature review on distributional finite elements. The distributional finite element de Rham complexes are adopted to construct equilibrated residual error estimators in [6], which are then extended to discrete distributional differential forms in [32], discrete distributional elasticity complexes in [17], and discrete distributional Hessian and divdiv complexes in [25] with applications in cohomology groups. Recently, in [16], the distributional finite element divdiv element has been constructed and applied for solving the mixed formulation of the biharmonic equation in arbitrary dimensions. The distributional finite elements allow the use of piecewise polynomials with less smoothness, which is especially useful for high-order differential operators.
Let us use a more familiar $2$nd order operator $\nabla^{2}$ as an example to illustrate the motivation. The $H^{2}$-conforming finite element on tetrahedron meshes [24, 13, 15, 43] requires polynomials of degree $9$ and above and possesses extra smoothness at vertices and edges. Therefore, it is hardly used in practice. Simple finite elements can be constructed if the differential operators are understood in the distribution sense.
For the discretization of the biharmonic equation in two dimensions, the so-called Hellan-Herrmann-Johnson (HHJ) mixed method [21, 22, 30] requires only normal-normal continuous finite elements for symmetric tensors and thus $C^{0}$-conforming Lagrange element, not $C^{1}$-conforming elements, can be used for displacement. This normal-normal continuous finite element is then employed to solve linear elasticity [36] and Reissner-Mindlin plates [37], and used to construct the first two-dimensional distributional finite element divdiv complexes in [11]. Recently, the distributional finite element divdiv element for solving the mixed formulation of the biharmonic equation has been extended to arbitrary dimensions in [16].
Now we move to the curl div operator. Introduce the space $H({\rm curl\,}\operatorname{div},\Omega;\mathbb{T}):=\left\{\boldsymbol{\tau}\in L^{2}(\Omega;\mathbb{T}):{\rm curl\,}\operatorname{div}\boldsymbol{\tau}\in L^{2}(\Omega;\mathbb{R}^{3})\right\}$, where $\mathbb{T}$ is the space of traceless tensors. A mixed formulation of the quad-curl problem is to find $\boldsymbol{\sigma}\in H(\operatorname{curl}\operatorname{div},\Omega;\mathbb{T})$, $\boldsymbol{u}\in L^{2}(\Omega;\mathbb{R}^{3})$ and $\phi\in H_{0}^{1}(\Omega)$ such that
$(\boldsymbol{\sigma},\boldsymbol{\tau})+b(\boldsymbol{\tau},\psi;\boldsymbol{u})$ $=0,$ $\boldsymbol{\tau}\in H(\operatorname{curl}\operatorname{div},\Omega;\mathbb{T}),\psi\in H_{0}^{1}(\Omega),$
$b(\boldsymbol{\sigma},\phi;\boldsymbol{v})$ $=-\langle\boldsymbol{f},\boldsymbol{v}\rangle,$ $\boldsymbol{v}\in L^{2}(\Omega;\mathbb{R}^{3}),$
where the bilinear form $b(\boldsymbol{\tau},\psi;\boldsymbol{v}):=(\operatorname{curl}{\operatorname{div}}\boldsymbol{\tau},\boldsymbol{v})+({\rm grad\,}\psi,\boldsymbol{v}).$ The term $({\rm grad\,}\psi,\boldsymbol{u})$ is introduced to impose the divergence free condition $\operatorname{div}\boldsymbol{u}=0$.
Finite element spaces conforming to $H({\rm curl\,}\operatorname{div},\Omega;\mathbb{T})$ are relatively complicated due to the smoothness requirement ${\rm curl\,}\operatorname{div}\boldsymbol{\tau}\in L^{2}(\Omega;\mathbb{R}^{3})$. In the distributional sense
$\langle{\rm curl\,}\operatorname{div}\boldsymbol{\tau},\boldsymbol{v}\rangle=-(\boldsymbol{\tau},{\rm grad\,}{\rm curl\,}\boldsymbol{v}),\quad\boldsymbol{v}\in H({\rm grad\,}{\rm curl\,},\Omega),$
the smoothness can be shifted to the test function $\boldsymbol{v}$, where $H({\rm grad\,}{\rm curl\,},\Omega):=\{\boldsymbol{u}\in L^{2}(\Omega;\mathbb{R}^{3}):{\rm curl\,}\boldsymbol{u}\in H^{1}(\Omega;\mathbb{R}^{3})\}$. Of course $H({\rm grad\,}{\rm curl\,})$-conforming finite elements are not easy to construct neither. For example, the $H({\rm grad\,}{\rm curl\,})$-conforming finite elements are constructed in [42, 26, 40, 14, 15], which requires polynomial of degree at least $7$ and dimension of shape function space at least $315$.
The key idea is to strike a balance of the smoothness of the trial function $\boldsymbol{\tau}$ and the test function $\boldsymbol{v}$. Given a mesh $\mathcal{T}_{h}$, let $H^{s}(\mathcal{T}_{h})$ be the space of piecewise $H^{s}$ function. Introduce the traceless tensor space with tangential-normal continuity
$\Sigma^{\rm tn}:=\{\boldsymbol{\tau}\in H^{1}(\mathcal{T}_{h};\mathbb{T}):\left.\llbracket\boldsymbol{n}\times\boldsymbol{\tau}\boldsymbol{n}\rrbracket\right|_{F}=0\text{ for each }F\in\mathring{\mathcal{F}}_{h}\},$
where $\llbracket\boldsymbol{n}\times\boldsymbol{\tau}\boldsymbol{n}\rrbracket$ is the jump of $\boldsymbol{n}\times\boldsymbol{\tau}\boldsymbol{n}$ across all interior faces $F$. While for the test space, we use space
$V_{0}^{{\rm curl\,}}:=H_{0}({\rm curl\,},\Omega)\cap H^{1}({\rm curl\,},\mathcal{T}_{h}),$
where $H^{1}({\rm curl\,},\mathcal{T}_{h}):=\{\boldsymbol{v}\in H^{1}(\mathcal{T}_{h};\mathbb{R}^{3}):{\rm curl\,}\boldsymbol{v}\in H^{1}(\mathcal{T}_{h};\mathbb{R}^{3})\}$. Define a weak operator $({\rm curl\,}\operatorname{div})_{w}:\Sigma^{\rm tn}\to(V_{0}^{{\rm curl\,}})^{\prime}$ by
(1) $\langle({\rm curl\,}\operatorname{div})_{w}\boldsymbol{\tau},\boldsymbol{v}\rangle:=\sum_{T\in\mathcal{T}_{h}}(\operatorname{div}\boldsymbol{\tau},{\rm curl\,}\boldsymbol{v})_{T}-\sum_{F\in\mathring{\mathcal{F}}_{h}}(\llbracket\boldsymbol{n}^{\intercal}\boldsymbol{\tau}\boldsymbol{n}\rrbracket,\boldsymbol{n}_{F}\cdot{\rm curl\,}\boldsymbol{v})_{F},$
which is analog to the weak divdiv operator in HHJ mixed method [21, 22, 30]. Now the function $\boldsymbol{\tau}$ is tangential-normally continuous and $\boldsymbol{v}$ is tangentially continuous so that $\boldsymbol{n}_{F}\cdot{\rm curl\,}\boldsymbol{v}=\operatorname*{rot}_{F}\boldsymbol{v}$ is continuous on face $F$. One can easily show $({\rm curl\,}\operatorname{div})_{w}\boldsymbol{\tau}={\rm curl\,}\operatorname{div}\boldsymbol{\tau}$ in the distribution sense by taking $\boldsymbol{v}\in C_{0}^{\infty}(\Omega;\mathbb{R}^{3})$ in (1).
We will use the tangential-normal continuous finite element constructed in [20] for the discretization of $\Sigma^{\rm tn}$. For an integer $k\geq 0$, take $\mathbb{P}_{k}(T;\mathbb{T})$ as the space of shape functions. The degrees of freedom (DoFs) are given by:
(2a) $\int_{F}\boldsymbol{t}_{i}^{\intercal}\boldsymbol{\tau}\boldsymbol{n}\,q\,{\rm d}S,\quad$ $q\in\mathbb{P}_{k}(F),i=1,2,F\in\mathcal{F}(T),$
(2b) $\int_{T}\boldsymbol{\tau}:\boldsymbol{q}\,{\rm d}x,\quad$ $\boldsymbol{q}\in\mathbb{P}_{k-1}(T;\mathbb{T}),$
where $\boldsymbol{t}_{1}$ and $\boldsymbol{t}_{2}$ denote two mutually perpendicular unit tangential vectors of face $F$ and are used to determine the tangential component of the vector $\boldsymbol{\tau}\boldsymbol{n}$. The global finite element space $\Sigma_{k,h}^{\rm tn}$ by requiring single valued (2a) is tangential-normally continuous.
We use Nédélec elements $\mathring{\mathbb{V}}_{h}^{{\rm curl\,}}\subset V_{0}^{{\rm curl\,}}$ for the tangential continuous vector space, and use the Riesz representation of the $L^{2}$-inner product to bring the abstract dual to a concrete function. Define $({\rm curl\,}\operatorname{div})_{h}:\Sigma_{h}^{\rm tn}\to\mathring{\mathbb{V}}_{h}^{{\rm curl\,}}$ such that
(3) $(({\rm curl\,}\operatorname{div})_{h}\boldsymbol{\tau}_{h},\boldsymbol{v}_{h})=\langle({\rm curl\,}\operatorname{div})_{w}\boldsymbol{\tau}_{h},\boldsymbol{v}_{h}\rangle,\quad\boldsymbol{v}_{h}\in\mathring{\mathbb{V}}_{h}^{{\rm curl\,}},$
and its $L^{2}$-adjoint operator $({\rm grad\,}{\rm curl\,})_{h}:\mathring{\mathbb{V}}_{h}^{{\rm curl\,}}\to\Sigma_{h}^{\rm tn}$.
By including the tensor version of the Nédélec elements $\mathbb{V}_{k,h}^{{\rm curl\,}}(\mathbb{M})$, and the Lagrange elements $\mathbb{V}_{k+1,h}^{{\rm grad\,}}(\mathbb{R}^{3})$, we are able to construct the distributional finite element ${\rm curl\,}\operatorname{div}$ complex:
(4) $\mathbb{R}^{3}\times\{0\}\rightarrow\mathbb{V}_{k+1,h}^{{\rm grad\,}}(\mathbb{R}^{3})\times\mathbb{R}$ $\xrightarrow{({\rm grad\,}\!,\,\operatorname{mskw}\boldsymbol{x})}\mathbb{V}_{k,h}^{{\rm curl\,}}(\mathbb{M})\xrightarrow{\operatorname{dev}{\rm curl\,}}$
$\Sigma_{k-1,h}^{\rm tn}\xrightarrow{({\rm curl\,}\operatorname{div})_{h}}\mathring{\mathbb{V}}_{(k,\ell),h}^{{\rm curl\,}}\xrightarrow{\operatorname{div}_{h}}\mathring{\mathbb{V}}_{\ell+1,h}^{{\rm grad\,}}\rightarrow 0,$
where $({\rm grad\,},\operatorname{mskw}\boldsymbol{x})\begin{pmatrix}\boldsymbol{v}\\ c\end{pmatrix}={\rm grad\,}\boldsymbol{v}+c\operatorname{mskw}\boldsymbol{x}$, and $\ell=k-1$ or $\ell=k$ is introduced to distinguish the first and second kind of Nédélec element. The lowest order, i.e. $k=1,\ell=0$, of the last three elements are illustrated in Fig. 1.
Figure 1. The simplest elements $\Sigma_{0,h}^{\rm tn}-\mathring{\mathbb{V}}_{(1,0),h}^{{\rm curl\,}}-\mathring{\mathbb{V}}_{1,h}^{{\rm grad\,}}$: the first is a piecewise constant traceless matrix with tangential-normal continuity, the second is the lowest order edge element, and the third is the linear Lagrange element.
The finite element complex (4) is a discretization of the distributional ${\rm curl\,}\operatorname{div}$ complex:
$\mathbb{R}^{3}\times\{0\}\rightarrow$ $H^{1}(\Omega;\mathbb{R}^{3})\times\mathbb{R}\xrightarrow{({\rm grad\,},\operatorname{mskw}\boldsymbol{x})}H(\operatorname{curl},\Omega;\mathbb{M})\xrightarrow{\mathrm{dev~{}curl}}$
$H^{-1}({\rm curl\,}\operatorname{div},\Omega;\mathbb{T})\xrightarrow{{\rm curl\,}\operatorname{div}}H^{-1}(\operatorname{div},\Omega)\xrightarrow{\operatorname{div}}H^{-1}(\Omega)\rightarrow 0,$
where $H^{-1}(\Omega):=(H_{0}^{1}(\Omega))^{\prime}$, and Sobolev spaces of negative order
$H^{-1}(\operatorname{div},\Omega)$ :=$:=\left\{\boldsymbol{u}\in H^{-1}(\Omega;\mathbb{R}^{3}):\operatorname{div}\boldsymbol{u}\in H^{-1}(\Omega)\right\},$
$H^{-1}({\rm curl\,}\operatorname{div},\Omega;\mathbb{T})$ :=$:=\left\{\boldsymbol{\tau}\in L^{2}(\Omega;\mathbb{T}):{\rm curl\,}\operatorname{div}\boldsymbol{\tau}\in H^{-1}(\operatorname{div},\Omega)\right\}.$
As an application, we consider the fourth-order curl problem $-{\rm curl\,}\Delta{\rm curl\,}\boldsymbol{u}=\boldsymbol{f},\operatorname{div}\boldsymbol{u}=0$ with boundary conditions $\boldsymbol{u}\times\boldsymbol{n}=\operatorname{curl}\boldsymbol{u}=0$ on $\partial\Omega$. The distributional mixed finite element method is to find $\boldsymbol{\sigma}_{h}\in{\Sigma}^{\rm tn}_{k-1,h}$, $\boldsymbol{u}_{h}\in\mathring{\mathbb{V}}_{(k,\ell),h}^{{\rm curl\,}}$ and $\phi_{h}\in\mathring{\mathbb{V}}_{\ell+1,h}^{{\rm grad\,}}$ such that
(5a) $(\boldsymbol{\sigma}_{h},\boldsymbol{\tau}_{h})+b_{h}(\boldsymbol{\tau}_{h},\psi_{h};\boldsymbol{u}_{h})$ $=0,$ $\boldsymbol{\tau}_{h}\in{\Sigma}^{\rm tn}_{k-1,h},\psi_{h}\in\mathring{\mathbb{V}}_{\ell+1,h}^{{\rm grad\,}},$
(5b) $b_{h}(\boldsymbol{\sigma}_{h},\phi_{h};\boldsymbol{v}_{h})$ $=-(\boldsymbol{f},\boldsymbol{v}_{h}),$ $\boldsymbol{v}_{h}\in\mathring{\mathbb{V}}_{(k,\ell),h}^{{\rm curl\,}},$
where $b_{h}(\boldsymbol{\tau},\psi;\boldsymbol{v}):=(({\rm curl\,}\operatorname{div})_{h}\boldsymbol{\tau},\boldsymbol{v})+({\rm grad\,}\psi,\boldsymbol{v})$ and $({\rm curl\,}\operatorname{div})_{h}$ is a discretization of distributional ${\rm curl\,}\operatorname{div}$ operator, cf. (3).
We prove two discrete inf-sup conditions and thus obtain the well-posedness of (5a)-(5b) and optimal order convergence
$\|\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h}\|_{0,h}+\|I_{h}^{{\rm curl\,}}\boldsymbol{u}-\boldsymbol{u}_{h}\|_{H(({\rm grad\,}{\rm curl\,})_{h})}$ $\lesssim h^{k}(|\boldsymbol{\sigma}|_{k}+|\boldsymbol{u}|_{k}),$
$\|\boldsymbol{u}-\boldsymbol{u}_{h}\|_{H({\rm curl\,})}+h\|\boldsymbol{u}-\boldsymbol{u}_{h}\|_{H(({\rm grad\,}{\rm curl\,})_{h})}$ $\lesssim h^{k}(|\boldsymbol{\sigma}|_{k}+|\boldsymbol{u}|_{k}+|{\rm curl\,}\boldsymbol{u}|_{k}).$
By the duality argument, the order of $\|{\rm curl\,}(I_{h}^{{\rm curl\,}}\boldsymbol{u}-\boldsymbol{u}_{h})\|$ can be improved to $h^{k+1}$ on convex domains. Both $\|I_{h}^{{\rm curl\,}}\boldsymbol{u}-\boldsymbol{u}_{h}\|_{H(({\rm grad\,}{\rm curl\,})_{h})}$ and $\|{\rm curl\,}(I_{h}^{{\rm curl\,}}\boldsymbol{u}-\boldsymbol{u}_{h})\|$ are superconvergent. Post-processing can be applied to improve the approximation to $\boldsymbol{u}$.
Furthermore, we apply hybridization techniques to (5a)-(5b), leading to a stabilization-free weak Galerkin method and extending to the $H({\rm grad\,}{\rm curl\,})$ nonconforming finite elements introduced in [28, 45] for solving the quad-curl problem. Equivalently, we identify the complex that accommodates these nonconforming finite elements and generalize them to arbitrary orders.
For other discretization of the quad-curl problem, we refer to the macro finite element method in [27], nonconforming finite element methods in [29, 39, 41], mixed finite element methods in [38, 10], decoupled finite element methods in [8, 44, 7], and references cited therein.
The rest of this paper is organized as follows. Section 2 focuses on the distributional ${\rm curl\,}\operatorname{div}$ complex. A distributional finite element ${\rm curl\,}\operatorname{div}$ complex is constructed in Section 3, and applied to solve the quad-curl problem in Section 4. The hybridization of the distributional mixed finite element method and the equivalence to other methods are presented in Section 5.
2. Distributional ${\rm curl\,}\operatorname{div}$ complex
In this section, we present the distributional ${\rm curl\,}\operatorname{div}$ complex and introduce the weak differential operator $({\rm curl\,}\operatorname{div})_{w}$ which can be defined on the tangential-normal continuous matrix functions.
2.1. Notation
Let $K\subset\mathbb{R}^{3}$ be a non-degenerated $3$-dimensional polyhedron. Denote by $\mathcal{F}(K)$ the set of all $2$-dimensional faces of $K$. For $F\in\mathcal{F}(K)$, denote by $\mathcal{E}(F)$ the set of all edges of $F$. Choose a normal vector $\boldsymbol{n}_{F}$ and two mutually perpendicular unit tangential vectors $\boldsymbol{t}_{F,1}$ and $\boldsymbol{t}_{F,2}$, which will be abbreviated as $\boldsymbol{t}_{1}$ and $\boldsymbol{t}_{2}$ for simplicity. Let $\boldsymbol{n}_{K}$ be the unit outward normal vector to $\partial K$, which will be abbreviated as $\boldsymbol{n}$. For $F\in\mathcal{F}(K)$ and $e\in\mathcal{E}(F)$, denote by $\boldsymbol{n}_{F,e}$ the unit vector being parallel to $F$ and outward normal to $\partial F$. Set $\boldsymbol{t}_{F,e}:=\boldsymbol{n}_{K}\times\boldsymbol{n}_{F,e}.$
Given a face $F\in\mathcal{F}(K)$, and a vector $\boldsymbol{v}\in\mathbb{R}^{3}$, define
$\Pi_{F}\boldsymbol{v}=(\boldsymbol{n}\times\boldsymbol{v})\times\boldsymbol{n}=(\boldsymbol{I}-\boldsymbol{n}\boldsymbol{n}^{\intercal})\boldsymbol{v}$
as the projection of $\boldsymbol{v}$ onto the face $F$ which is called the tangential component of $\boldsymbol{v}$. The vector $\boldsymbol{n}\times\boldsymbol{v}=(\boldsymbol{n}\times\Pi_{F})\boldsymbol{v}$ is called the tangential trace of $\boldsymbol{v}$, which is a rotation of $\Pi_{F}\boldsymbol{v}$ on $F$ ($90^{\circ}$ counter-clockwise with respect to $\boldsymbol{n}$).
Define the surface gradient operator as $\nabla_{F}:=\Pi_{F}\nabla$. For a scalar function $v$, define the surface ${\rm curl\,}$:
${\rm curl\,}_{F}v=\boldsymbol{n}\times\nabla v=\boldsymbol{n}\times\nabla_{F}v.$
For a vector function $\boldsymbol{v}$, the surface rot operator is defined as:
${\rm rot}_{F}\boldsymbol{v}:=(\boldsymbol{n}\times\nabla)\cdot\boldsymbol{v}=(\boldsymbol{n}\times\nabla_{F})\cdot\Pi_{F}\boldsymbol{v}=\boldsymbol{n}\cdot({\rm curl\,}\boldsymbol{v}),$
which represents the normal component of ${\rm curl\,}\boldsymbol{v}$.
Denote the space of all $3\times 3$ matrices by $\mathbb{M}$, and all trace-free/traceless $3\times 3$ matrices by $\mathbb{T}$. Define the deviation $\operatorname{dev}\boldsymbol{\tau}=\boldsymbol{\tau}-\frac{1}{3}(\operatorname{tr}\boldsymbol{\tau})\boldsymbol{I}$. Obviously for a scalar function $u$, $\operatorname{dev}(u\boldsymbol{I})=0$. For a vector $\boldsymbol{w}=(\omega_{1},\omega_{2},\omega_{3})^{\intercal}\in\mathbb{R}^{3}$, let
$\operatorname{mskw}\boldsymbol{w}:=\begin{pmatrix}0&-\omega_{3}&\omega_{2}\\ \omega_{3}&0&-\omega_{1}\\ -\omega_{2}&\omega_{1}&0\end{pmatrix}.$
For a tensor-valued function $\boldsymbol{\tau}$, $\operatorname{div}\boldsymbol{\tau}$ and ${\rm curl\,}\boldsymbol{\tau}$ mean operators $\operatorname{div}$ and ${\rm curl\,}$ are applied row-wisely to $\boldsymbol{\tau}$. By direct calculation, we have the identities:
(6) $\operatorname{div}\operatorname{mskw}\boldsymbol{v}=-{\rm curl\,}\boldsymbol{v},\quad(\operatorname{mskw}\boldsymbol{v})\boldsymbol{n}=\boldsymbol{v}\times\boldsymbol{n}.$
We use $\{\mathcal{T}_{h}\}_{h>0}$ to denote a shape regular family of simplicial meshes of $\Omega$ with mesh size $h=\max_{T\in\mathcal{T}_{h}}h_{T}$ and $h_{T}$ being the diameter of $T$. Let $\mathcal{F}_{h}$, ̊$\mathring{\mathcal{F}}_{h}$, $\mathcal{E}_{h}$, ̊$\mathring{\mathcal{E}}_{h}$, $\mathcal{V}_{h}$ and $\mathring{\mathcal{V}}_{h}$ be the set of all faces, interior faces, edges, interior edges, vertices and interior vertices of $\mathcal{T}_{h}$, respectively. Let $T\in\mathcal{T}_{h}$ be a tetrahedron with four vertices $\texttt{v}_{0},\ldots,\texttt{v}_{3}$. Denote by $\lambda_{i}$ the $i$th barycentric coordinate with respect to the simplex $T$ for $i=0,\ldots,3$. Set $\boldsymbol{t}_{ij}:=\texttt{v}_{j}-\texttt{v}_{i}$ as the edge vector from $\texttt{v}_{i}$ to $\texttt{v}_{j}$.
Given a nonnegative integer $k$, let $\mathbb{P}_{k}(T)$ stand for the set of all polynomials in $T$ with the total degree no more than $k$, and let $\mathbb{P}_{k}(T;\mathbb{X})$ denote the tensor or vector version with $\mathbb{X}=\mathbb{R}^{3}$, $\mathbb{M}$ and $\mathbb{T}$. When $k<0$, set $\mathbb{P}_{k}(T):=\{0\}.$
[…]
Introduce the following Sobolev spaces
$H({\rm curl\,},D)$ :=$:=\{\boldsymbol{u}\in L^{2}(D;\mathbb{R}^{3}):{\rm curl\,}\boldsymbol{u}\in L^{2}(D;\mathbb{R}^{3})\},$
$H(\operatorname{div},D)$ :=$:=\{\boldsymbol{u}\in L^{2}(D;\mathbb{R}^{3}):\operatorname{div}\boldsymbol{u}\in L^{2}(D)\},$
$H({\rm grad\,}{\rm curl\,},D)$ :=$:=\{\boldsymbol{u}\in L^{2}(D;\mathbb{R}^{3}):{\rm curl\,}\boldsymbol{u}\in H^{1}(D;\mathbb{R}^{3})\},$
$H(\operatorname{curl}\operatorname{div},D;\mathbb{T})$ :=$:=\left\{\boldsymbol{\tau}\in L^{2}(D;\mathbb{T}):\operatorname{curl}\operatorname{div}\boldsymbol{\tau}\in L^{2}(D;\mathbb{R}^{3})\right\},$
where $H^{s}(D;\mathbb{X}):=H^{s}(D)\otimes\mathbb{X}$. Define piecewise smooth function space, for $s>0$,
$H^{s}(\mathcal{T}_{h}):=\{v\in L^{2}(\Omega):v\mid_{T}\in H^{s}(T)\text{ for all }T\in\mathcal{T}_{h}\},$
and $H^{s}(\mathcal{T}_{h};\mathbb{X})$ its tensor or vector version with $\mathbb{X}=\mathbb{R}^{3}$, $\mathbb{M}$ and $\mathbb{T}$. Let ${\rm grad\,}_{\mathcal{T}_{h}}$ and ${\rm curl\,}_{\mathcal{T}_{h}}$ be the elementwise version of ${\rm grad\,}$ and ${\rm curl\,}$ associated with $\mathcal{T}_{h}$, respectively.
2.2. The ${\rm curl\,}\operatorname{div}$ complexes
The ${\rm curl\,}\operatorname{div}$ complex in three dimensions reads as [3, (47)]
[…]
$H({\rm curl\,}\operatorname{div},\Omega;\mathbb{T})\xrightarrow{{\rm curl\,}\operatorname{div}}H(\operatorname{div},\Omega)\xrightarrow{\operatorname{div}}L^{2}(\Omega)\rightarrow 0.$
When $\Omega$ is topologically trivial, i.e., all co-homology group of $\Omega$ is trivial, then (7) is exact. The smoothness of the potential can be further improved to be in $H^{1}$.
It is difficult to construct $H(\operatorname{curl}\operatorname{div},\Omega;\mathbb{T})$-conforming finite element with lower order degree of polynomials. To relax the smoothness, we are going to present a distributional ${\rm curl\,}\operatorname{div}$ complex with negative Sobolev spaces involved.
Define
$H^{-1}(\operatorname{div},\Omega)=\{\boldsymbol{v}\in H^{-1}(\Omega;\mathbb{R}^{3}):\operatorname{div}\boldsymbol{v}\in H^{-1}(\Omega)\}.$
In [12], we have shown that $(H_{0}({\rm curl\,},\Omega))^{\prime}=H^{-1}(\operatorname{div},\Omega).$
Define
$H^{-1}(\operatorname{curl}\operatorname{div},\Omega;\mathbb{T}):=\left\{\boldsymbol{\tau}\in L^{2}(\Omega;\mathbb{T}):\operatorname{curl}\operatorname{div}\boldsymbol{\tau}\in H^{-1}(\operatorname{div},\Omega)\right\}$
with squared norm
$\|\boldsymbol{\tau}\|_{H(\operatorname{curl}\operatorname{div})}^{2}:=\|\boldsymbol{\tau}\|^{2}+\|{\rm curl\,}\operatorname{div}\boldsymbol{\tau}\|_{H^{-1}(\operatorname{div})}^{2}=\|\boldsymbol{\tau}\|^{2}+\|{\rm curl\,}\operatorname{div}\boldsymbol{\tau}\|_{-1}^{2}.$
Lemma 2.1.
The distributional ${\rm curl\,}\operatorname{div}$ complex in three dimensions is
(8) $\mathbb{R}^{3}\times\{0\}\rightarrow$ $H^{1}(\Omega;\mathbb{R}^{3})\times\mathbb{R}\xrightarrow{({\rm grad\,},\operatorname{mskw}\boldsymbol{x})}H(\operatorname{curl},\Omega;\mathbb{M})\xrightarrow{\mathrm{dev~{}curl}}$
$H^{-1}({\rm curl\,}\operatorname{div},\Omega;\mathbb{T})\xrightarrow{{\rm curl\,}\operatorname{div}}H^{-1}(\operatorname{div},\Omega)\xrightarrow{\operatorname{div}}H^{-1}(\Omega)\rightarrow 0.$
When $\Omega\subset\mathbb{R}^{3}$ is a bounded and topologically trivial Lipschitz domain, (8) is exact.
Proof.
Apparently (8) is a complex. The surjection $\operatorname{div}H^{-1}(\operatorname{div},\Omega)=H^{-1}(\Omega)$ follows from $\operatorname{div}L^{2}(\Omega;\mathbb{R}^{3})=H^{-1}(\Omega)$ and $\operatorname{div}L^{2}(\Omega;\mathbb{R}^{3})\subseteq\operatorname{div}H^{-1}(\operatorname{div},\Omega)$. We then verify its exactness.
[…]
For $\boldsymbol{v}\in H^{-1}(\operatorname{div},\Omega)\cap\ker(\operatorname{div})$, by the exactness of the de Rham complex [18], there exists $\boldsymbol{\tau}\in H^{1}(\Omega;\mathbb{M})$ such that $\boldsymbol{v}={\rm curl\,}\operatorname{div}\boldsymbol{\tau}$. Notice that ${\rm curl\,}\operatorname{div}(p\boldsymbol{I})={\rm curl\,}{\rm grad\,}p=0$. Then $\boldsymbol{v}={\rm curl\,}\operatorname{div}(\operatorname{dev}\boldsymbol{\tau})\in{\rm curl\,}\operatorname{div}H^{-1}({\rm curl\,}\operatorname{div},\Omega;\mathbb{T})$.
[…]
For $\boldsymbol{\tau}\in H^{-1}({\rm curl\,}\operatorname{div},\Omega;\mathbb{T})\cap\ker({\rm curl\,}\operatorname{div})$, by the de Rham complex, there exists a function $u\in L^{2}(\Omega)$ s.t. $\operatorname{div}\boldsymbol{\tau}={\rm grad\,}u$. Then $\operatorname{div}(\boldsymbol{\tau}-u\boldsymbol{I})=0$, which means $\boldsymbol{\tau}=u\boldsymbol{I}+{\rm curl\,}\boldsymbol{\sigma}$ with $\boldsymbol{\sigma}\in H^{1}(\Omega;\mathbb{M})$. By the traceless of $\boldsymbol{\tau}$, we get $\boldsymbol{\tau}=\operatorname{dev}{\rm curl\,}\boldsymbol{\sigma}\in\operatorname{dev}{\rm curl\,}H^{1}(\Omega;\mathbb{M})\subseteq\operatorname{dev}{\rm curl\,}H(\operatorname{curl},\Omega;\mathbb{M})$.
[…]
Since ${\rm curl\,}(\operatorname{mskw}\boldsymbol{x})=2\boldsymbol{I}$, we have ${\rm grad\,}H^{1}(\Omega;\mathbb{R}^{3})\cap\textrm{span}\{\operatorname{mskw}\boldsymbol{x}\}=\{0\}$. For $\boldsymbol{\tau}\in H(\operatorname{curl},\Omega;\mathbb{M})\cap\ker(\operatorname{dev}{\rm curl\,})$, we have ${\rm curl\,}\boldsymbol{\tau}=\frac{1}{3}\operatorname{tr}({\rm curl\,}\boldsymbol{\tau})\boldsymbol{I}$. Apply $\operatorname{div}$ on both sides to get ${\rm grad\,}(\operatorname{tr}({\rm curl\,}\boldsymbol{\tau}))=0$. Then $\operatorname{tr}({\rm curl\,}\boldsymbol{\tau})$ is constant, and ${\rm curl\,}\boldsymbol{\tau}=2c\boldsymbol{I}$ with $c\in\mathbb{R}$. This implies ${\rm curl\,}(\boldsymbol{\tau}-c\operatorname{mskw}\boldsymbol{x})=0$. Therefore $\boldsymbol{\tau}\in{\rm grad\,}H^{1}(\Omega;\mathbb{R}^{3})\oplus\textrm{span}\{\operatorname{mskw}\boldsymbol{x}\}$. ∎
Next we use the framework developed in [12] to present a Helmholtz decomposition of $H^{-1}({\rm curl\,}\operatorname{div},\Omega;\mathbb{T})$. Denote by
$K^{c}=H_{0}({\rm curl\,},\Omega)\cap\ker(\operatorname{div})=\{\boldsymbol{v}\in H_{0}({\rm curl\,},\Omega):\boldsymbol{v}\perp{\rm grad\,}H_{0}^{1}(\Omega)\}.$
Then ${\rm curl\,}{\rm curl\,}:K^{c}\to H^{-1}(\operatorname{div},\Omega)\cap\ker(\operatorname{div})$ is isomorphic. Indeed, by the Helmholtz decomposition $L^{2}(\Omega;\mathbb{R}^{3})={\rm curl\,}K^{c}+\nabla H^{1}(\Omega)$ [2], we have
${\rm curl\,}{\rm curl\,}K^{c}={\rm curl\,}L^{2}(\Omega;\mathbb{R}^{3})=H^{-1}(\operatorname{div},\Omega)\cap\ker(\operatorname{div}).$
Lemma 2.2.
It holds the Helmholtz decomposition
$H^{-1}({\rm curl\,}\operatorname{div},\Omega;\mathbb{T})=\operatorname{dev}{\rm curl\,}H({\rm curl\,},\Omega;\mathbb{M})\oplus\operatorname{mskw}K^{c}.$
Proof.
With complex (8) and identity (6), we build up the commutative diagram
[…]
Apply the framework in [12] to get the required Helmholtz decomposition. ∎
We introduce the tensor space with tangential-normal continuity
$\Sigma^{\rm tn}:=\{\boldsymbol{\tau}\in H^{1}(\mathcal{T}_{h};\mathbb{T}):\left.\llbracket\boldsymbol{n}\times\boldsymbol{\tau}\boldsymbol{n}\rrbracket\right|_{F}=0\text{ for each }F\in\mathring{\mathcal{F}}_{h}\},$
where $\llbracket\boldsymbol{n}\times\boldsymbol{\tau}\boldsymbol{n}\rrbracket$ is the jump of $\boldsymbol{n}\times\boldsymbol{\tau}\boldsymbol{n}$ across $F$. Let space $V_{0}^{{\rm curl\,}}:=H_{0}({\rm curl\,},\Omega)\cap H^{1}({\rm curl\,},\mathcal{T}_{h}),$ where $H^{1}({\rm curl\,},\mathcal{T}_{h}):=\{\boldsymbol{v}\in H^{1}(\mathcal{T}_{h};\mathbb{R}^{3}):{\rm curl\,}\boldsymbol{v}\in H^{1}(\mathcal{T}_{h};\mathbb{R}^{3})\}$. We define a weak operator $({\rm curl\,}\operatorname{div})_{w}:\Sigma^{\rm tn}\to(V_{0}^{{\rm curl\,}})^{\prime}\subset(C_{0}^{\infty}(\Omega;\mathbb{R}^{3}))^{\prime}$ by
(9) $\langle({\rm curl\,}\operatorname{div})_{w}\boldsymbol{\tau},\boldsymbol{v}\rangle:=\sum_{T\in\mathcal{T}_{h}}(\operatorname{div}\boldsymbol{\tau},{\rm curl\,}\boldsymbol{v})_{T}-\sum_{F\in\mathring{\mathcal{F}}_{h}}(\llbracket\boldsymbol{n}^{\intercal}\boldsymbol{\tau}\boldsymbol{n}\rrbracket,\boldsymbol{n}_{F}\cdot{\rm curl\,}\boldsymbol{v})_{F}.$
Notice that only interior faces are included in the second term as $\boldsymbol{n}_{F}\cdot{\rm curl\,}\boldsymbol{v}=\operatorname*{rot}_{F}\boldsymbol{v}=0$ for $\boldsymbol{v}\in H_{0}({\rm curl\,},\Omega)$.
Lemma 2.3.
For $\boldsymbol{\tau}\in\Sigma^{\rm tn}$, the following identity holds in the distribution sense
$({\rm curl\,}\operatorname{div})_{w}\boldsymbol{\tau}={\rm curl\,}\operatorname{div}\boldsymbol{\tau}.$
Proof.
By the definition of the distributional derivative and employing the integration by parts element-wise we get for $\boldsymbol{v}\in C_{0}^{\infty}(\Omega;\mathbb{R}^{3})$ that
$\langle{\rm curl\,}\operatorname{div}\boldsymbol{\tau},\boldsymbol{v}\rangle$ $:=-(\boldsymbol{\tau},{\rm grad\,}{\rm curl\,}\boldsymbol{v})=\sum_{T\in\mathcal{T}_{h}}(\operatorname{div}\boldsymbol{\tau},{\rm curl\,}\boldsymbol{v})_{T}-\sum_{T\in\mathcal{T}_{h}}(\boldsymbol{\tau}\boldsymbol{n},{\rm curl\,}\boldsymbol{v})_{\partial T}$
$\;=\sum_{T\in\mathcal{T}_{h}}(\operatorname{div}\boldsymbol{\tau},{\rm curl\,}\boldsymbol{v})_{T}-\sum_{T\in\mathcal{T}_{h}}(\boldsymbol{n}^{\intercal}\boldsymbol{\tau}\boldsymbol{n},\boldsymbol{n}\cdot{\rm curl\,}\boldsymbol{v})_{\partial T}$
$\;\quad-\sum_{T\in\mathcal{T}_{h}}(\boldsymbol{n}\times\boldsymbol{\tau}\boldsymbol{n},\boldsymbol{n}\times{\rm curl\,}\boldsymbol{v})_{\partial T}.$
As $\boldsymbol{n}\times\boldsymbol{\tau}\boldsymbol{n}$ is continuous and $\boldsymbol{v}\in C_{0}^{\infty}(\Omega;\mathbb{R}^{3})$, the last term is canceled. Then rearrange the second term face-wisely to derive
$\langle{\rm curl\,}\operatorname{div}\boldsymbol{\tau},\boldsymbol{v}\rangle=\langle({\rm curl\,}\operatorname{div})_{w}\boldsymbol{\tau},\boldsymbol{v}\rangle,\quad\forall~{}\boldsymbol{v}\in C_{0}^{\infty}(\Omega;\mathbb{R}^{3}).$
Thus, $({\rm curl\,}\operatorname{div})_{w}\boldsymbol{\tau}={\rm curl\,}\operatorname{div}\boldsymbol{\tau}$ in the distribution sense. ∎
Similarly we can define the weak operator $({\rm grad\,}{\rm curl\,})_{w}:V_{0}^{{\rm curl\,}}\to(\Sigma^{\rm tn})^{\prime}$ as
$\langle({\rm grad\,}{\rm curl\,})_{w}\boldsymbol{v},\boldsymbol{\tau}\rangle:=\sum_{T\in\mathcal{T}_{h}}(\boldsymbol{\tau},{\rm grad\,}{\rm curl\,}\boldsymbol{v})_{T}-\sum_{F\in\mathcal{F}_{h}}(\boldsymbol{n}\times\boldsymbol{\tau}\boldsymbol{n},\llbracket\boldsymbol{n}\times{\rm curl\,}\boldsymbol{v}\rrbracket)_{F}.$
By definition, we have the duality
(10) $\langle({\rm curl\,}\operatorname{div})_{w}\boldsymbol{\tau},\boldsymbol{v}\rangle=-\langle\boldsymbol{\tau},({\rm grad\,}{\rm curl\,})_{w}\boldsymbol{v}\rangle,\quad\boldsymbol{\tau}\in\Sigma^{\rm tn},\boldsymbol{v}\in V_{0}^{{\rm curl\,}}.$
When $\boldsymbol{\tau}\in H({\rm curl\,}\operatorname{div},\Omega;\mathbb{T})\cap\Sigma^{\rm tn}$, $\langle({\rm curl\,}\operatorname{div})_{w}\boldsymbol{\tau},\boldsymbol{v}\rangle=({\rm curl\,}\operatorname{div}\boldsymbol{\tau},\boldsymbol{v})$ and when $\boldsymbol{v}\in H({\rm grad\,}{\rm curl\,},\Omega)$, $\langle\boldsymbol{\tau},({\rm grad\,}{\rm curl\,})_{w}\boldsymbol{v}\rangle=(\boldsymbol{\tau},{\rm grad\,}{\rm curl\,}\boldsymbol{v})$. The duality (10) strikes a balance of the smoothness of $\boldsymbol{\tau}$ and $\boldsymbol{v}$ so that the second order differential operators can be defined for less smooth functions.
3. Distributional finite element ${\rm curl\,}\operatorname{div}$ complex
We shall construct a finite element counterpart of the distributional ${\rm curl\,}\operatorname{div}$ complex (8).
3.1. Finite element spaces
We first recall the tangential-normal continuous finite element for traceless tensors in [20]. Take $\mathbb{P}_{k}(T;\mathbb{T})$ as the space of shape functions with $k\geq 0$. The degrees of freedom (DoFs) are given by
(11a) $\int_{F}\boldsymbol{t}_{i}^{\intercal}\boldsymbol{\tau}\boldsymbol{n}\,q\,{\rm d}S,\quad$ $q\in\mathbb{P}_{k}(F),i=1,2,F\in\mathcal{F}(T),$
(11b) $\int_{T}\boldsymbol{\tau}:\boldsymbol{q}\,{\rm d}x,\quad$ $\boldsymbol{q}\in\mathbb{P}_{k-1}(T;\mathbb{T}).$
In order to give a geometric decomposition of space $\mathbb{P}_{k}(T;\mathbb{T})$, we present two intrinsic bases of $\mathbb{T}$ which are variants of a basis constructed in [23].
(a) Basis $\{\operatorname{dev}(\nabla\lambda_{i}\otimes\boldsymbol{t}_{i\ell}),\operatorname{dev}(\nabla\lambda_{j}\otimes\boldsymbol{t}_{j\ell})\}_{\ell=0}^{3}$.
(b) Basis $\{\boldsymbol{t}_{mi}\otimes\nabla\lambda_{\ell},\boldsymbol{t}_{mj}\otimes\nabla\lambda_{\ell}\}_{\ell=0}^{3}$.
Figure 2. Two intrinsic bases of traceless matrix $\mathbb{T}$.
Lemma 3.1.
Let $(ij\ell m)$ be a cyclic permutation of $(0123)$. Then the set
(12) $\{\operatorname{dev}(\nabla\lambda_{i}\otimes\boldsymbol{t}_{i\ell}),\operatorname{dev}(\nabla\lambda_{j}\otimes\boldsymbol{t}_{j\ell}),\ell=0,\ldots,3\}$
is dual to
(13) $\{\boldsymbol{t}_{mi}\otimes\nabla\lambda_{\ell},\boldsymbol{t}_{mj}\otimes\nabla\lambda_{\ell}\}_{\ell=0}^{3}.$
Consequently both are bases of $\mathbb{T}$.
Proof.
The duality follows from the identity $\boldsymbol{t}_{ij}\cdot\nabla\lambda_{\ell}=\delta_{j\ell}-\delta_{i\ell}$, where $\delta_{i\ell}$ is the Kronecker delta function, and $\boldsymbol{t}\otimes\nabla\lambda_{\ell}:\boldsymbol{I}=\boldsymbol{t}\cdot\nabla\lambda_{\ell}=0$ for vector $\boldsymbol{t}$ tangent to $F_{\ell}$. ∎
As $\nabla\lambda_{i}\parallel\boldsymbol{n}_{F_{i}}$, the basis (13) is face-wise and each face contributes two while the basis (12) is vertex-wise. They are illustrated in Fig. 2.
Define
$\operatorname{tr}_{F}^{\rm tn}\boldsymbol{\tau}=\boldsymbol{n}_{F}\times\boldsymbol{\tau}\boldsymbol{n}_{F},$
and $\operatorname{tr}^{\rm tn}:C(T;\mathbb{T})\to L^{2}(\partial T;\mathbb{R}^{2})$ as $\operatorname{tr}^{\rm tn}|_{F}=\operatorname{tr}_{F}^{\rm tn}$. Let the bubble polynomial space of degree $k$
$\mathbb{B}_{k}^{\rm tn}(T;\mathbb{T}):=\{\boldsymbol{\tau}\in\mathbb{P}_{k}(T;\mathbb{T}):\operatorname{tr}^{\rm tn}\boldsymbol{\tau}=0\}.$
Notice that for the identity matrix $\boldsymbol{I}$, $\operatorname{tr}^{\rm tn}\boldsymbol{I}=0$ but $\boldsymbol{I}\not\in\mathbb{B}_{k}^{\rm tn}(T;\mathbb{T})$ as ${\rm trace}(\boldsymbol{I})\neq 0$.
Lemma 3.2.
For $\ell,i=0,\ldots,3,i\neq\ell$, we have
$\lambda_{\ell}\operatorname{dev}(\nabla\lambda_{i}\otimes\boldsymbol{t}_{i\ell})\in\mathbb{B}_{k}^{\rm tn}(T;\mathbb{T}).$
Proof.
Let $(ij\ell m)$ be a cyclic permutation of $(0123)$. The edge $\texttt{v}_{i}\texttt{v}_{\ell}$ is contained in faces $F_{m}$ and $F_{j}$ and thus $\boldsymbol{t}_{i\ell}\cdot\boldsymbol{n}_{F}=0$ for $F=F_{m},F_{j}$. The vector $\nabla\lambda_{i}\parallel\boldsymbol{n}_{F_{i}}$ and thus $\boldsymbol{n}_{F_{i}}\times\nabla\lambda_{i}=0$ on $F_{i}$. On the face $F_{\ell}$, $\lambda_{\ell}|_{F_{\ell}}=0$. Notice that the identity matrix $\boldsymbol{I}$ satisfies $\operatorname{tr}^{\rm tn}\boldsymbol{I}=0$.
So we have verified $\operatorname{tr}^{\rm tn}(\lambda_{\ell}\operatorname{dev}(\nabla\lambda_{i}\otimes\boldsymbol{t}_{i\ell}))=0$. ∎
By changing $\nabla\lambda_{i}$ to the parallel vector $\boldsymbol{n}_{i}$, we present the following geometric decomposition of $\mathbb{P}_{k}(T;\mathbb{T})$.
Lemma 3.3.
We have the geometric decomposition
[…]
where we give a characterization of the bubble space
$\mathbb{B}_{k}^{\rm tn}(T;\mathbb{T})=\mathbb{P}_{k-1}(T)\otimes\mathrm{span}\{\lambda_{\ell}\operatorname{dev}(\boldsymbol{n}_{i}\otimes\boldsymbol{t}_{i\ell}),\lambda_{\ell}\operatorname{dev}(\boldsymbol{n}_{j}\otimes\boldsymbol{t}_{j\ell}),\ell=0,\ldots,3\}.$
Define the global finite element space
${\Sigma}_{k,h}^{\rm tn}:=\{\boldsymbol{\tau}_{h}\in L^{2}(\Omega;\mathbb{T})$ $:\boldsymbol{\tau}_{h}|_{T}\in\mathbb{P}_{k}(T;\mathbb{T})\textrm{ for each }T\in\mathcal{T}_{h},$
[…]
Remark 3.4.
The construction can be readily extended to arbitrary dimension $\mathbb{R}^{d}$ for $d\geq 2$. There are $d+1$ faces for a $d$-simplex. At each face, we have $d-1$ linearly independent traceless matrices $\{\boldsymbol{t}_{F,i}\otimes\boldsymbol{n}_{F}\}_{i=1}^{d-1}$ and in total $(d+1)(d-1)$ such matrices form a basis of $\mathbb{T}$. The basis of bubble functions is constructed vertices-wisely $\{\lambda_{\ell}\operatorname{dev}(\nabla\lambda_{i}\otimes\boldsymbol{t}_{i\ell})\}$ for $\ell=0,\ldots,d$ and $d-1$ different $i$ for each $\ell$. ∎
Remark 3.5.
Since the basis functions are expressed in terms of intrinsic geometric quantities—tangential and normal vectors along with barycentric coordinates—the defined finite element spaces are affine invariant, meaning the space remains unchanged under affine transformations. Thus, the traditional method of using a reference tetrahedron and transforming to the physical element can be applied, resulting in the same finite element space. In particular, curved elements can be handled using the Piola transform, as described in equations (3.76)-(3.77) of [33]. Extending the two-dimensional work of [4] to a distributional mixed finite element method specifically tailored for the quad-curl problem on domains with curved boundaries is an interesting topic for future research. ∎
As $H^{-1}(\operatorname{div},\Omega)=(H_{0}({\rm curl\,},\Omega))^{\prime}$, we can use $H({\rm curl\,})$-conforming finite elements, i.e. Nédélec elements [34, 35] for the pair space. Take
$\mathcal{N}_{k,\ell}^{c}(T):=\boldsymbol{x}\times\mathbb{P}_{k-1}(T;\mathbb{R}^{3})+{\rm grad\,}\mathbb{P}_{\ell+1}(T),\quad\ell=k\text{ or }k-1$
as the space of shape functions. The element $\mathcal{N}_{k,k-1}^{c}$ is the first kind and $\mathcal{N}_{k,k}^{c}$ is the second kind Nédélec elements [34, 35]. The degrees of freedom $\mathcal{N}_{k,\ell}^{c}(T)$ are given by
(14a) $(\boldsymbol{v}\cdot\boldsymbol{t},q)_{e},$ $\quad q\in\mathbb{P}_{\ell}(e),e\in\mathcal{E}(T),$
(14b) $(\boldsymbol{n}\times\boldsymbol{v}\times\boldsymbol{n},\boldsymbol{q})_{F},$ $\quad\boldsymbol{q}\in{\rm curl\,}_{F}\mathbb{P}_{k-1}(F)\oplus\boldsymbol{x}\mathbb{P}_{\ell-2}(F),F\in\mathcal{F}(T),$
(14c) $(\boldsymbol{v},\boldsymbol{q})_{T},$ $\quad\boldsymbol{q}\in{\rm curl\,}\mathbb{P}_{k-2}(T;\mathbb{R}^{3})\oplus\boldsymbol{x}\mathbb{P}_{\ell-3}(T).$
Define global finite element spaces
$\mathbb{V}_{(k,\ell),h}^{{\rm curl\,}}:=\{\boldsymbol{v}_{h}\in H({\rm curl\,},\Omega):$ $\boldsymbol{v}_{h}|_{T}\in\mathcal{N}_{k,\ell}^{c}(T)\textrm{ for}T\in\mathcal{T}_{h}$
[…]
Let $\mathring{\mathbb{V}}_{(k,\ell),h}^{{\rm curl\,}}:=\mathbb{V}_{(k,\ell),h}^{{\rm curl\,}}\cap H_{0}({\rm curl\,},\Omega)$ and $\mathbb{V}_{k,h}^{{\rm curl\,}}:=\mathbb{V}_{(k,k),h}^{{\rm curl\,}}$.
We use the standard Lagrange element for $H^{1}(\Omega)$
$\mathbb{V}_{\ell+1,h}^{{\rm grad\,}}:=\{\psi_{h}\in H^{1}(\Omega):\psi_{h}|_{T}\in\mathbb{P}_{\ell+1}(T)\textrm{ for }T\in\mathcal{T}_{h}\},$
and let $\mathring{\mathbb{V}}_{\ell+1,h}^{{\rm grad\,}}:=\mathbb{V}_{\ell+1,h}^{{\rm grad\,}}\cap H_{0}^{1}(\Omega)$. Let
$\mathbb{V}_{k+1,h}^{{\rm grad\,}}(\mathbb{R}^{3}):=\mathbb{V}_{k+1,h}^{{\rm grad\,}}\otimes\mathbb{R}^{3},\quad\text{ and }\quad\mathbb{V}_{k,h}^{{\rm curl\,}}(\mathbb{M}):=\mathbb{R}^{3}\otimes\mathbb{V}_{k,h}^{{\rm curl\,}}.$
The degree of polynomial may be skipped in the notation of finite element spaces when it is clear from the context.
3.2. Distributional finite element complex
By treating the right hand side of (9) as a bilinear form defined on $\boldsymbol{\Sigma}^{\rm tn}\times V_{0}^{{\rm curl\,}}$, the weak operators can be naturally extended to the discrete spaces by restricting the bilinear form to subspaces. Define
$({\rm curl\,}\operatorname{div})_{h}=J_{h}({\rm curl\,}\operatorname{div})_{w}:{\Sigma}_{h}^{\rm tn}\to(\mathring{\mathbb{V}}_{h}^{{\rm curl\,}})^{\prime}\cong\mathring{\mathbb{V}}_{h}^{{\rm curl\,}},$
$({\rm grad\,}{\rm curl\,})_{h}=J_{h}({\rm grad\,}{\rm curl\,})_{w}:\mathring{\mathbb{V}}_{h}^{{\rm curl\,}}\to({\Sigma}_{h}^{\rm tn})^{\prime}\cong{\Sigma}_{h}^{\rm tn},$
where the isomorphism $J_{h}$ is the Reisz representation of the $L^{2}$-inner product and realized by the inverse of the mass matrix. More precisely, for $\boldsymbol{\sigma}_{h}\in\Sigma_{h}^{\rm tn}$, $({\rm curl\,}\operatorname{div})_{h}\boldsymbol{\sigma}_{h}\in\mathring{\mathbb{V}}_{h}^{{\rm curl\,}}$ such that
$(({\rm curl\,}\operatorname{div})_{h}\boldsymbol{\sigma}_{h},\boldsymbol{v}_{h})$ $=\langle({\rm curl\,}\operatorname{div})_{w}\boldsymbol{\sigma}_{h},\boldsymbol{v}_{h}\rangle$
(15) $=\sum_{T\in\mathcal{T}_{h}}(\operatorname{div}\boldsymbol{\sigma}_{h},{\rm curl\,}\boldsymbol{v}_{h})_{T}-\sum_{F\in\mathring{\mathcal{F}}_{h}}(\llbracket\boldsymbol{n}^{\intercal}\boldsymbol{\sigma}_{h}\boldsymbol{n}\rrbracket,\boldsymbol{n}_{F}\cdot{\rm curl\,}\boldsymbol{v}_{h})_{F}$
$=-\sum_{T\in\mathcal{T}_{h}}(\boldsymbol{\sigma}_{h},{\rm grad\,}{\rm curl\,}\boldsymbol{v}_{h})_{T}+\sum_{F\in\mathcal{F}_{h}}(\boldsymbol{n}\times\boldsymbol{\sigma}_{h}\boldsymbol{n},\llbracket\boldsymbol{n}\times{\rm curl\,}\boldsymbol{v}_{h}\rrbracket)_{F}$
$=-(\boldsymbol{\sigma}_{h},({\rm grad\,}{\rm curl\,})_{h}\boldsymbol{v}_{h}).$
For a fixed triangulation $\mathcal{T}_{h}$, $({\rm curl\,}\operatorname{div})_{h}$ is well defined which can be obtained by inverting the mass matrix of $\mathring{\mathbb{V}}_{h}^{{\rm curl\,}}$. However $\{({\rm curl\,}\operatorname{div})_{h}\}$ is not uniformly bounded when $h\to 0$ as (15) is not a bounded linear functional of $L^{2}(\Omega)$.
Similarly define discrete div operator $\operatorname{div}_{h}:H_{0}({\rm curl\,},\Omega)\to\mathring{\mathbb{V}}_{h}^{{\rm grad\,}}$ by
$(\operatorname{div}_{h}\boldsymbol{v},\psi_{h})=-(\boldsymbol{v},{\rm grad\,}\psi_{h}),\quad\psi_{h}\in\mathring{\mathbb{V}}_{h}^{{\rm grad\,}}.$
Notice that $\operatorname{div}_{h}$ is the $L^{2}$-adjoint of $-{\rm grad\,}$ restricted to $\mathring{\mathbb{V}}_{h}^{{\rm grad\,}}$. Again for a fixed triangulation $\mathcal{T}_{h}$, $\operatorname{div}_{h}$ is well defined which can be obtained by inverting the mass matrix of $\mathring{\mathbb{V}}_{h}^{{\rm grad\,}}$. However $\{\operatorname{div}_{h}\}$ is not uniformly bounded as $h\to 0$ as $(\boldsymbol{v},{\rm grad\,}\psi)$ is in general not a bounded linear functional of $L^{2}$ unless $\boldsymbol{v}\in H(\operatorname{div},\Omega)$.
Theorem 3.6.
The distributional finite element ${\rm curl\,}\operatorname{div}$ complex is
(16) $\mathbb{R}^{3}\times\{0\}\rightarrow\mathbb{V}_{k+1,h}^{{\rm grad\,}}(\mathbb{R}^{3})\times\mathbb{R}$ $\xrightarrow{({\rm grad\,}\!,\,\operatorname{mskw}\boldsymbol{x})}\mathbb{V}_{k,h}^{{\rm curl\,}}(\mathbb{M})\xrightarrow{\operatorname{dev}{\rm curl\,}}$
$\Sigma_{k-1,h}^{\rm tn}$ $\xrightarrow{({\rm curl\,}\operatorname{div})_{h}}\mathring{\mathbb{V}}_{(k,\ell),h}^{{\rm curl\,}}\xrightarrow{\operatorname{div}_{h}}\mathring{\mathbb{V}}_{\ell+1,h}^{{\rm grad\,}}\rightarrow 0.$
When $\Omega\subset\mathbb{R}^{3}$ is a bounded and topologically trivial Lipschitz domain, (16) is exact.
Proof.
As ${\rm curl\,}{\rm grad\,}=0$, it is straightforward to verify $\operatorname{div}_{h}({\rm curl\,}\operatorname{div})_{h}=0$. Take $\boldsymbol{\tau}_{h}=\operatorname{dev}{\rm curl\,}\boldsymbol{\sigma}_{h}$ with $\boldsymbol{\sigma}_{h}\in\mathbb{V}_{h}^{{\rm curl\,}}(\mathbb{M})$. Since $\boldsymbol{n}\times\boldsymbol{\tau}_{h}\boldsymbol{n}=\boldsymbol{n}\times({\rm curl\,}\boldsymbol{\sigma}_{h})\boldsymbol{n}$ is single-value across each face $F\in\mathring{\mathcal{F}}_{h}$, we get $\boldsymbol{\tau}_{h}\in{\Sigma}^{\rm tn}_{h}$. In the distribution sense, ${\rm curl\,}\operatorname{div}\operatorname{dev}{\rm curl\,}=0$ and so is $({\rm curl\,}\operatorname{div})_{h}\operatorname{dev}{\rm curl\,}$. In summary, we have verified (16) is a complex.
Verification of $\ker({\rm grad\,},\operatorname{mskw}\boldsymbol{x})=\mathbb{R}^{3}\times\{0\}$ is trivial. For $\boldsymbol{\tau}_{h}\in\mathbb{V}_{k,h}^{{\rm curl\,}}(\mathbb{M})$ and $\operatorname{dev}{\rm curl\,}(\boldsymbol{\tau}_{h})=0$, by the exactness of (8), we can find $(\boldsymbol{v},c)\in H^{1}(\Omega;\mathbb{R}^{3})\times\mathbb{R}$ s.t. ${\rm grad\,}\boldsymbol{v}+c\operatorname{mskw}\boldsymbol{x}=\boldsymbol{\tau}_{h}$. As $\boldsymbol{\tau}_{h}$ is polynomial of degree $k$, we conclude $\boldsymbol{v}\in\mathbb{V}_{k+1,h}^{{\rm grad\,}}(\mathbb{R}^{3})$.
From the finite element de Rham complex ${\rm grad\,}\mathring{\mathbb{V}}_{\ell+1,h}^{{\rm grad\,}}=\mathring{\mathbb{V}}_{(k,\ell),h}^{{\rm curl\,}}\cap\ker({\rm curl\,})$, we have
(17) $\operatorname{div}_{h}\mathring{\mathbb{V}}_{(k,\ell),h}^{{\rm curl\,}}=\mathring{\mathbb{V}}_{\ell+1,h}^{{\rm grad\,}}.$
It remains to prove
(18) $\operatorname{dev}{\rm curl\,}\mathbb{V}_{k,h}^{{\rm curl\,}}(\mathbb{M})$ $=\Sigma_{k-1,h}^{\rm tn}\cap\ker(({\rm curl\,}\operatorname{div})_{h}),$
(19) $({\rm curl\,}\operatorname{div})_{h}\Sigma_{k-1,h}^{\rm tn}$ $=\mathring{\mathbb{V}}_{(k,\ell),h}^{{\rm curl\,}}\cap\ker(\operatorname{div}_{h}).$
We will prove (19) in Corollary 3.12 and (18) in Corollary 3.13. ∎
3.3. Characterization of null spaces
Define $K_{h}^{c}=\mathring{\mathbb{V}}_{(k,\ell),h}^{{\rm curl\,}}\cap\ker(\operatorname{div}_{h})$ and $({\rm curl\,}{\rm curl\,})_{h}:K_{h}^{c}\to K_{h}^{c}$ so that
(20) $(({\rm curl\,}{\rm curl\,})_{h}\boldsymbol{u}_{h},\boldsymbol{v}_{h})=({\rm curl\,}\boldsymbol{u}_{h},{\rm curl\,}\boldsymbol{v}_{h}),\quad\boldsymbol{v}_{h}\in K_{h}^{c}.$
Lemma 3.7.
We have the discrete Poincaré inequalities
(21) $\|\boldsymbol{v}_{h}\|$ $\lesssim\|{\rm curl\,}\boldsymbol{v}_{h}\|,\quad\quad\quad\;\;\;\,\boldsymbol{v}_{h}\in K_{h}^{c},$
(22) $\|\boldsymbol{v}_{h}\|_{H({\rm curl\,})}$ $\lesssim\|({\rm curl\,}{\rm curl\,})_{h}\boldsymbol{v}_{h}\|,\quad\boldsymbol{v}_{h}\in K_{h}^{c}.$
Proof.
The first Poincaré inequality (21) can be found in [19, Lemma 3.4 and Theorem 3.6] and [33, Lemma 7.20]. Consequently $({\rm curl\,}\cdot,{\rm curl\,}\cdot)$ is an inner product on $K_{h}^{c}$ and the operator $({\rm curl\,}{\rm curl\,})_{h}$ is isomorphic.
Taking $\boldsymbol{u}_{h}=\boldsymbol{v}_{h}$ in (20) and applying (21), we get
$\|{\rm curl\,}\boldsymbol{v}_{h}\|^{2}\leq\|({\rm curl\,}{\rm curl\,})_{h}\boldsymbol{v}_{h}\|\|\boldsymbol{v}_{h}\|\lesssim\|({\rm curl\,}{\rm curl\,})_{h}\boldsymbol{v}_{h}\|\|{\rm curl\,}\boldsymbol{v}_{h}\|.$
Hence $\|{\rm curl\,}\boldsymbol{v}_{h}\|\lesssim\|({\rm curl\,}{\rm curl\,})_{h}\boldsymbol{v}_{h}\|.$ The proof is finished by applying (21) again to bound $\|\boldsymbol{v}_{h}\|$. ∎
We will introduce interpolation operators satisfying the following commutative diagrams:
Figure 3. Identities connecting the weak differential operators and interpolation operators.
Let $I_{h}^{\rm tn}:\Sigma^{\rm tn}\to{\Sigma}^{\rm tn}_{k-1,h}$ be the interpolation operator using DoF (11) and denote by $\boldsymbol{\sigma}_{I}=I_{h}^{\rm tn}\boldsymbol{\sigma}$.
Lemma 3.8.
For $\boldsymbol{\tau}\in\Sigma^{\rm tn}$ and $\boldsymbol{v}_{h}\in\mathring{\mathbb{V}}_{(k,\ell),h}^{{\rm curl\,}}$, it holds
(23) $(({\rm curl\,}\operatorname{div})_{h}(I_{h}^{\rm tn}\boldsymbol{\tau}),\boldsymbol{v}_{h})=\langle({\rm curl\,}\operatorname{div})_{w}\boldsymbol{\tau},\boldsymbol{v}_{h}\rangle.$
Proof.
By definition,
$\sum_{T\in\mathcal{T}_{h}}-(\boldsymbol{\tau}-I_{h}^{\rm tn}\boldsymbol{\tau},{\rm grad\,}{\rm curl\,}\boldsymbol{v}_{h})_{T}+(\boldsymbol{n}\times(\boldsymbol{\tau}-I_{h}^{\rm tn}\boldsymbol{\tau})\boldsymbol{n},\boldsymbol{n}\times{\rm curl\,}\boldsymbol{v}_{h})_{\partial T}=0,$
which is true due to DoFs (11a)-(11b) and the fact ${\rm grad\,}{\rm curl\,}\boldsymbol{v}_{h}|_{T}\in\mathbb{P}_{k-2}(T;\mathbb{R}^{3})$ and $\boldsymbol{n}\times{\rm curl\,}\boldsymbol{v}_{h}|_{F}\in\mathbb{P}_{k-1}(F;\mathbb{R}^{2})$. ∎
We introduce the interpolation operator $I_{h}^{{\rm curl\,}}:V_{0}^{{\rm curl\,}}\to\mathring{\mathbb{V}}_{(k,\ell),h}^{{\rm curl\,}}$ defined by DoF (14) and denote by $\boldsymbol{v}_{I}=I_{h}^{{\rm curl\,}}\boldsymbol{v}$.
Lemma 3.9.
For $\boldsymbol{v}\in V_{0}^{{\rm curl\,}}$ and $\boldsymbol{\tau}_{h}\in\Sigma^{\rm tn}_{k-1,h}$, it holds
(24) $\langle({\rm curl\,}\operatorname{div})_{w}\boldsymbol{\tau}_{h},\boldsymbol{v}\rangle=(({\rm curl\,}\operatorname{div})_{h}\boldsymbol{\tau}_{h},I_{h}^{{\rm curl\,}}\boldsymbol{v}).$
Proof.
By integration by parts and $\boldsymbol{\tau}_{h}\mid_{T}\in\mathbb{P}_{k-1}(T;\mathbb{T})$, we have
$\quad\langle({\rm curl\,}\operatorname{div})_{w}\boldsymbol{\tau}_{h},\boldsymbol{v}\rangle$
$=$ $\sum_{T\in\mathcal{T}_{h}}({\rm curl\,}\operatorname{div}\boldsymbol{\tau}_{h},\boldsymbol{v})_{T}-\sum_{F\in\mathring{\mathcal{F}}_{h}}([\boldsymbol{n}\times\operatorname{div}\boldsymbol{\tau}_{h}+{\rm curl\,}_{F}(\boldsymbol{n}^{\intercal}\boldsymbol{\tau}_{h}\boldsymbol{n})],\boldsymbol{v})_{F}$
$-\sum_{e\in\mathring{\mathcal{E}}_{h}}([\boldsymbol{n}^{\intercal}\boldsymbol{\tau}_{h}\boldsymbol{n}]_{e},\boldsymbol{v}\cdot\boldsymbol{t}_{e})_{e}$
$=$ $\sum_{T\in\mathcal{T}_{h}}({\rm curl\,}\operatorname{div}\boldsymbol{\tau}_{h},\boldsymbol{v}_{I})_{T}-\sum_{F\in\mathring{\mathcal{F}}_{h}}([\boldsymbol{n}\times\operatorname{div}\boldsymbol{\tau}_{h}+{\rm curl\,}_{F}(\boldsymbol{n}^{\intercal}\boldsymbol{\tau}_{h}\boldsymbol{n})],\boldsymbol{v}_{I})_{F}$
$-\sum_{e\in\mathring{\mathcal{E}}_{h}}([\boldsymbol{n}^{\intercal}\boldsymbol{\tau}_{h}\boldsymbol{n}]_{e},\boldsymbol{v}_{I}\cdot\boldsymbol{t}_{e})_{e}=(({\rm curl\,}\operatorname{div})_{h}\boldsymbol{\tau}_{h},\boldsymbol{v}_{I}),$
where $[\boldsymbol{n}^{\intercal}\boldsymbol{\tau}_{h}\boldsymbol{n}]_{e}=\sum_{T\in\mathcal{T}_{h}}\sum_{F\in\mathcal{F}(T),e\subset\partial F}(\boldsymbol{n}^{\intercal}\boldsymbol{\tau}_{h}|_{T}\boldsymbol{n})|_{e}(\boldsymbol{t}_{e}\cdot\boldsymbol{t}_{F,e}).$
Corollary 3.10.
Let $\boldsymbol{\tau}_{h}\in{\Sigma}^{\rm tn}_{k-1,h}$ satisfy $({\rm curl\,}\operatorname{div})_{h}\boldsymbol{\tau}_{h}=0$. Then we have ${\rm curl\,}\operatorname{div}\boldsymbol{\tau}_{h}=0$ and $\boldsymbol{\tau}_{h}\in H(\operatorname{curl}\operatorname{div},\Omega;\mathbb{T})$.
Proof.
Let $\boldsymbol{\tau}_{h}\in{\Sigma}^{\rm tn}_{h}$ satisfy $({\rm curl\,}\operatorname{div})_{h}\boldsymbol{\tau}_{h}=0$. By Lemma 3.9 and Lemma 2.3, ${\rm curl\,}\operatorname{div}\boldsymbol{\tau}_{h}=0$ in the distribution sense. ∎
At first glance, the null space $\Sigma_{k-1,h}^{\rm tn}\cap\ker(({\rm curl\,}\operatorname{div})_{h})$ is larger than the null space $\Sigma_{k-1,h}^{\rm tn}\cap\ker({\rm curl\,}\operatorname{div})$ as the test function is $\mathring{\mathbb{V}}_{(k,\ell),h}^{{\rm curl\,}}$ not $C_{0}^{\infty}$. However Lemma 3.9 implies that they are the same due to the design of finite element spaces and weak differential operators. This is in the same spirit of Hellan-Herrmann-Johnson (HHJ) element [21, 22, 30, 11] in 2D.
Lemma 3.11.
It holds
(25) $({\rm curl\,}\operatorname{div})_{h}(I_{h}^{\rm tn}(\operatorname{mskw}\boldsymbol{u}_{h}))=-({\rm curl\,}{\rm curl\,})_{h}\boldsymbol{u}_{h},\quad\boldsymbol{u}_{h}\in K_{h}^{c}.$
Proof.
Since $(\operatorname{mskw}\boldsymbol{u}_{h})\boldsymbol{n}=\boldsymbol{u}_{h}\times\boldsymbol{n}$, it follows that $\operatorname{mskw}\boldsymbol{u}_{h}\in\Sigma^{\rm tn}$. By the fact $\operatorname{div}\operatorname{mskw}\boldsymbol{u}_{h}=-{\rm curl\,}\boldsymbol{u}_{h}$, we have
$\langle({\rm curl\,}\operatorname{div})_{w}(\operatorname{mskw}\boldsymbol{u}_{h}),\boldsymbol{v_{h}}\rangle=(\operatorname{div}\operatorname{mskw}\boldsymbol{u}_{h},{\rm curl\,}\boldsymbol{v_{h}})=-({\rm curl\,}\boldsymbol{u}_{h},{\rm curl\,}\boldsymbol{v_{h}}).$
Then the result holds from (23). ∎
With complex (16) and identity (25), we have the commutative diagram
[…]
Corollary 3.12.
It holds
(27) $({\rm curl\,}\operatorname{div})_{h}\Sigma_{k-1,h}^{\rm tn}=K_{h}^{c}=\mathring{\mathbb{V}}_{(k,\ell),h}^{{\rm curl\,}}\cap\ker(\operatorname{div}_{h}).$
Proof.
It is straightforward to verify that $({\rm curl\,}\operatorname{div})_{h}{\Sigma}^{\rm tn}_{h}\subseteq K_{h}^{c}$. On the other side, take $\boldsymbol{w}_{h}\in K_{h}^{c}$. Let $\boldsymbol{u}_{h}=({\rm curl\,}{\rm curl\,})_{h}^{-1}\boldsymbol{w}_{h}\in K_{h}^{c}$ and set $\boldsymbol{\tau}_{h}=-I_{h}^{\rm tn}(\operatorname{mskw}\boldsymbol{u}_{h})\in{\Sigma}^{\rm tn}_{h}$. By (25),
$({\rm curl\,}\operatorname{div})_{h}\boldsymbol{\tau}_{h}=({\rm curl\,}{\rm curl\,})_{h}\boldsymbol{u}_{h}=\boldsymbol{w}_{h},$
which ends the proof. ∎
Corollary 3.13.
We have
(28) $\operatorname{dev}{\rm curl\,}\mathbb{V}_{k,h}^{{\rm curl\,}}(\mathbb{M})=\Sigma_{k-1,h}^{\rm tn}\cap\ker(({\rm curl\,}\operatorname{div})_{h}).$
Proof.
By complex (16), $\operatorname{dev}{\rm curl\,}\mathbb{V}_{k,h}^{{\rm curl\,}}(\mathbb{M})\subseteq\Sigma_{k-1,h}^{\rm tn}\cap\ker(({\rm curl\,}\operatorname{div})_{h})$. Then we prove (28) by dimension count. By (17) and (27),
$\dim\Sigma_{k-1,h}^{\rm tn}\cap\ker(({\rm curl\,}\operatorname{div})_{h})$ $=\dim\Sigma_{k-1,h}^{\rm tn}-\dim\mathring{\mathbb{V}}_{(k,k-1),h}^{{\rm curl\,}}+\dim\mathring{\mathbb{V}}_{k,h}^{{\rm grad\,}}$
$=|\mathring{\mathcal{V}}_{h}|-|\mathring{\mathcal{E}}_{h}|+2{k+1\choose 2}|{\mathcal{F}}_{h}|-{k+1\choose 2}|\mathring{\mathcal{F}}_{h}|+|\mathring{\mathcal{F}}_{h}|$
$\quad+|\mathcal{T}_{h}|\bigg{(}8{k+1\choose 3}-3{k\choose 3}+{k-1\choose 3}\bigg{)}.$
Similarly,
$\dim\operatorname{dev}{\rm curl\,}\mathbb{V}_{k,h}^{{\rm curl\,}}(\mathbb{M})$ $=\dim\mathbb{V}_{k,h}^{{\rm curl\,}}(\mathbb{M})-\dim\mathbb{V}_{k+1,h}^{{\rm grad\,}}(\mathbb{R}^{3})+2$
$=-3|{\mathcal{V}}_{h}|+3|{\mathcal{E}}_{h}|+3{k+1\choose 2}|{\mathcal{F}}_{h}|-3|{\mathcal{F}}_{h}|$
$\quad+|{\mathcal{T}}_{h}|\bigg{(}9{k\choose 3}-3{k-1\choose 3}\bigg{)}+2.$
Combine the last two identities to get
$\quad\dim\Sigma_{k-1,h}^{\rm tn}\cap\ker(({\rm curl\,}\operatorname{div})_{h})-\dim\operatorname{dev}{\rm curl\,}\mathbb{V}_{k,h}^{{\rm curl\,}}(\mathbb{M})$
$=|\mathring{\mathcal{V}}_{h}|+3|{\mathcal{V}}_{h}|-|\mathring{\mathcal{E}}_{h}|-3|{\mathcal{E}}_{h}|+|\mathring{\mathcal{F}}_{h}|+3|{\mathcal{F}}_{h}|-4|{\mathcal{T}}_{h}|-2$
+($\quad+{k+1\choose 2}(4|{\mathcal{T}}_{h}|-|{\mathcal{F}}_{h}|-|\mathring{\mathcal{F}}_{h}|).$
Finally, we conclude the result from $4|{\mathcal{T}}_{h}|=|{\mathcal{F}}_{h}|+|\mathring{\mathcal{F}}_{h}|$, and the Euler’s formulas $|{\mathcal{V}}_{h}|-|{\mathcal{E}}_{h}|+|{\mathcal{F}}_{h}|-|{\mathcal{T}}_{h}|=1$ and $|\mathring{\mathcal{V}}_{h}|-|\mathring{\mathcal{E}}_{h}|+|\mathring{\mathcal{F}}_{h}|-|{\mathcal{T}}_{h}|=-1$. ∎
Remark 3.14.
Here is another proof of (28). Take $\boldsymbol{\tau}_{h}\in\Sigma^{\rm tn}_{k-1,h}\cap\ker(\!({\rm curl\,}\operatorname{div})_{h})$. By Corollary 3.10, we have ${\rm curl\,}\operatorname{div}\boldsymbol{\tau}_{h}=0$. Therefore $\boldsymbol{\tau}_{h}=\operatorname{dev}{\rm curl\,}\boldsymbol{\sigma}$ with $\boldsymbol{\sigma}\in H^{1}(\Omega;\mathbb{M})$. Let $\boldsymbol{\sigma}_{I}\in\mathbb{V}_{h}^{{\rm curl\,}}(\mathbb{M})$ be the interpolation of $\boldsymbol{\sigma}$ based on DoFs (14a)-(14c). The edge moment is not well defined for $H^{1}$ function but may be fixed by the fact $\operatorname{dev}{\rm curl\,}\boldsymbol{\sigma}=\boldsymbol{\tau}_{h}$ has extra smoothness. It is easy to verify that all the DoFs (11a)-(11b) of $\boldsymbol{\tau}_{h}-\operatorname{dev}{\rm curl\,}\boldsymbol{\sigma}_{I}$ vanish. Therefore $\boldsymbol{\tau}_{h}=\operatorname{dev}{\rm curl\,}\boldsymbol{\sigma}_{I}\in\operatorname{dev}{\rm curl\,}\mathbb{V}_{h}^{{\rm curl\,}}(\mathbb{M})$.
3.4. Helmholtz decompositions
The right half of the distributional finite element ${\rm curl\,}\operatorname{div}$ complex (16) is listed below
$\mathbb{V}_{k,h}^{{\rm curl\,}}(\mathbb{M})\xrightarrow{\operatorname{dev}{\rm curl\,}}\Sigma_{k-1,h}^{\rm tn}\xrightarrow{({\rm curl\,}\operatorname{div})_{h}}\mathring{\mathbb{V}}_{(k,\ell),h}^{{\rm curl\,}}\xrightarrow{\operatorname{div}_{h}}\mathring{\mathbb{V}}_{\ell+1,h}^{{\rm grad\,}}\rightarrow 0,\quad\ell=k\text{ or }k-1.$
By taking the dual, we have the short exact sequence
$0\leftarrow({\rm grad\,}{\rm curl\,})_{h}K_{h}^{c}\xleftarrow{({\rm grad\,}{\rm curl\,})_{h}}\mathring{\mathbb{V}}_{(k,\ell),h}^{{\rm curl\,}}=K_{h}^{c}\oplus{\rm grad\,}(\mathring{\mathbb{V}}_{\ell+1,h}^{{\rm grad\,}})\xleftarrow{{\rm grad\,}}\mathring{\mathbb{V}}_{\ell+1,h}^{{\rm grad\,}}\leftarrow 0$
By the framework in [12], we get the following Helmholtz decompositions from the last two complexes, commutative diagram (26) and (25).
Corollary 3.15.
We have the discrete Helmholtz decompositions
$\Sigma_{k-1,h}^{\rm tn}$ $=\operatorname{dev}{\rm curl\,}\mathbb{V}_{k,h}^{{\rm curl\,}}(\mathbb{M})\oplus^{L^{2}}({\rm grad\,}{\rm curl\,})_{h}K_{h}^{c},$
$\Sigma_{k-1,h}^{\rm tn}$ $=\operatorname{dev}{\rm curl\,}\mathbb{V}_{k,h}^{{\rm curl\,}}(\mathbb{M})\oplus I_{h}^{\rm tn}(\operatorname{mskw}K_{h}^{c}).$
Corollary 3.16.
We have the $L^{2}$-orthogonal Helmholtz decomposition of space $\mathring{\mathbb{V}}_{(k,\ell),h}^{{\rm curl\,}}$
(29) $\mathring{\mathbb{V}}_{(k,\ell),h}^{{\rm curl\,}}$ $=K_{h}^{c}\oplus^{L^{2}}{\rm grad\,}\mathring{\mathbb{V}}_{\ell+1,h}^{{\rm grad\,}}=({\rm curl\,}{\rm curl\,})_{h}K_{h}^{c}\oplus^{L^{2}}{\rm grad\,}\mathring{\mathbb{V}}_{\ell+1,h}^{{\rm grad\,}}$
$=({\rm curl\,}\operatorname{div})_{h}\Sigma_{k-1,h}^{\rm tn}\oplus^{L^{2}}{\rm grad\,}\mathring{\mathbb{V}}_{\ell+1,h}^{{\rm grad\,}}.$
Lemma 3.17.
We have the discrete Poincaré inequality
(30) $\|\boldsymbol{v}_{h}\|_{H({\rm curl\,})}\lesssim\|({\rm grad\,}{\rm curl\,})_{h}\boldsymbol{v}_{h}\|,\quad\boldsymbol{v}_{h}\in K_{h}^{c}.$
Proof.
Set $\boldsymbol{\tau}_{h}=I_{h}^{\rm tn}(\operatorname{mskw}\boldsymbol{v}_{h})\in\Sigma_{k-1,h}^{\rm tn}$. Then $({\rm curl\,}\operatorname{div})_{h}\boldsymbol{\tau}_{h}=-({\rm curl\,}{\rm curl\,})_{h}\boldsymbol{v}_{h}$ follows from (25). By the scaling argument, the inverse inequality and the Poincaré inequality (21),
(31) $\|\boldsymbol{\tau}_{h}\|\lesssim\|\boldsymbol{v}_{h}\|\lesssim\|{\rm curl\,}\boldsymbol{v}_{h}\|.$
It follows that
$\|{\rm curl\,}\boldsymbol{v}_{h}\|^{2}=(({\rm curl\,}{\rm curl\,})_{h}\boldsymbol{v}_{h},\boldsymbol{v}_{h})=-(({\rm curl\,}\operatorname{div})_{h}\boldsymbol{\tau}_{h},\boldsymbol{v}_{h})=(\boldsymbol{\tau}_{h},({\rm grad\,}{\rm curl\,})_{h}\boldsymbol{v}_{h}).$
Applying Cauchy-Schwarz inequality and (31), we obtain
$\|{\rm curl\,}\boldsymbol{v}_{h}\|\lesssim\|({\rm grad\,}{\rm curl\,})_{h}\boldsymbol{v}_{h}\|,$
which implies (30). ∎
4. Mixed finite element method of the quad-curl problem
Let $\Omega\subset\mathbb{R}^{3}$ be a bounded polygonal domain. Consider the fourth order problem
(32) {$-{\rm curl\,}\Delta{\rm curl\,}\boldsymbol{u}&=\boldsymbol{f}&&\text{ in }\Omega,\\ \operatorname{div}\boldsymbol{u}&=0&&\text{ in }\Omega,\\ \boldsymbol{u}\times\boldsymbol{n}=\operatorname{curl}\boldsymbol{u}\times\boldsymbol{n}&=0&&\text{ on }\partial\Omega,\end{aligned}\right.$
where $\boldsymbol{f}\in H^{-1}(\operatorname{div},\Omega)\cap\ker(\operatorname{div})$ is known. Such problem arises from multiphysics simulation such as modeling a magnetized plasma in magnetohydrodynamics [31, 9].
4.1. Distributional mixed formulation
Introducing $\boldsymbol{\sigma}:={\rm grad\,}{\rm curl\,}\boldsymbol{u}$, we have $\operatorname{tr}\boldsymbol{\sigma}=\operatorname{tr}{\rm grad\,}{\rm curl\,}\boldsymbol{u}=\operatorname{div}{\rm curl\,}\boldsymbol{u}=0$. Then rewrite problem (32) as the second-order system
(33) $\boldsymbol{\sigma}-{\rm grad\,}{\rm curl\,}\boldsymbol{u}&=0\quad\quad\text{ in }\Omega,\\ {\rm curl\,}{\operatorname{div}}\boldsymbol{\sigma}&=-\boldsymbol{f}\quad\text{ in }\Omega,\\ \operatorname{div}\boldsymbol{u}&=0\qquad\text{ in }\Omega,\\ \boldsymbol{u}\times\boldsymbol{n}=\operatorname{curl}\boldsymbol{u}&=0\quad\;\;\;\text{ on }\partial\Omega.\end{aligned}\right.$
A mixed formulation of the system (33) is to find $\boldsymbol{\sigma}\in H^{-1}(\operatorname{curl}\operatorname{div},\Omega;\mathbb{T})$, $\boldsymbol{u}\in H_{0}({\rm curl\,},\Omega)$ and $\phi\in H_{0}^{1}(\Omega)$ such that
(34a) $(\boldsymbol{\sigma},\boldsymbol{\tau})+b(\boldsymbol{\tau},\psi;\boldsymbol{u})$ $=0,$ $\boldsymbol{\tau}\in H^{-1}(\operatorname{curl}\operatorname{div},\Omega;\mathbb{T}),\psi\in H_{0}^{1}(\Omega),$
(34b) $b(\boldsymbol{\sigma},\phi;\boldsymbol{v})$ $=-\langle\boldsymbol{f},\boldsymbol{v}\rangle,$ $\boldsymbol{v}\in H_{0}({\rm curl\,},\Omega),$
where the bilinear form $b(\cdot,\cdot;\cdot):(H^{-1}(\operatorname{curl}\operatorname{div},\Omega;\mathbb{T})\times H_{0}^{1}(\Omega))\times H_{0}({\rm curl\,},\Omega)$
$b(\boldsymbol{\tau},\psi;\boldsymbol{v}):=\langle\operatorname{curl}{\operatorname{div}}\boldsymbol{\tau},\boldsymbol{v}\rangle+({\rm grad\,}\psi,\boldsymbol{v}).$
The term $({\rm grad\,}\psi,\boldsymbol{u})$ is introduced to impose the divergence free condition $\operatorname{div}\boldsymbol{u}=0$.
Lemma 4.1.
For $\boldsymbol{v}\in H_{0}({\rm curl\,},\Omega)$, it holds
(35) $\|\boldsymbol{v}\|_{H({\rm curl\,})}\lesssim\sup_{\boldsymbol{\tau}\in H^{-1}(\operatorname{curl}\operatorname{div},\Omega;\mathbb{T}),\psi\in H_{0}^{1}(\Omega)}\frac{b(\boldsymbol{\tau},\psi;\boldsymbol{v})}{\|\boldsymbol{\tau}\|_{H^{-1}(\operatorname{curl}\operatorname{div})}+|\psi|_{1}}.$
Proof of this lemma is similar to, indeed simpler than, that of the discrete inf-sup condition cf. Lemma 4.4, we thus skip the details here.
Lemma 4.2.
For $\boldsymbol{\tau}\in H^{-1}(\operatorname{curl}\operatorname{div},\Omega;\mathbb{T})$ and $\psi\in H_{0}^{1}(\Omega)$ satisfying
(36) $b(\boldsymbol{\tau},\psi;\boldsymbol{v})=0,\quad\forall\ \boldsymbol{v}\in H_{0}({\rm curl\,},\Omega),$
it holds
(37) $\|\boldsymbol{\tau}\|_{H^{-1}(\operatorname{curl}\operatorname{div})}^{2}+|\psi|_{1}^{2}=\|\boldsymbol{\tau}\|^{2}.$
Proof.
By taking $\boldsymbol{v}={\rm grad\,}\psi$ in (36), we get $\psi=0$. Then (36) becomes
$\langle\operatorname{curl}{\operatorname{div}}\boldsymbol{\tau},\boldsymbol{v}\rangle=0,\quad\forall\ \boldsymbol{v}\in H_{0}({\rm curl\,},\Omega).$
Hence $\operatorname{curl}{\operatorname{div}}\boldsymbol{\tau}=0$, and (37) follows. ∎
Combining (35), (37) and the Babuška-Brezzi theory [5] yileds the wellposedness of the mixed formulation (34a)-(34b).
Theorem 4.3.
The mixed formulation (34a)-(34b) is well-posed. Namely for any $\boldsymbol{f}\in H^{-1}(\operatorname{div},\Omega)\cap\ker(\operatorname{div})$, there exists a unique solution $(\boldsymbol{\sigma},\boldsymbol{u},\phi)$ to (34a)-(34b). Furthermore we have $\phi=0$, and the stability
(38) $\|\boldsymbol{\sigma}\|_{H^{-1}(\operatorname{curl}\operatorname{div})}+\|\boldsymbol{u}\|_{H({\rm curl\,})}\lesssim\|\boldsymbol{f}\|_{H^{-1}(\operatorname{div})}.$
Proof.
Combine the inf-sup condition (35) and the coercivity (37) to get (38) and the well-posedness of the mixed formulation (34a)-(34b). By choosing $\boldsymbol{v}={\rm grad\,}\phi$ in (34b), we get $\phi=0$ from $\operatorname{div}\boldsymbol{f}=0$. ∎
4.2. Distributional mixed finite element method
For $(\boldsymbol{\tau},\psi)\in\Sigma^{\rm tn}\times H_{0}^{1}(\Omega)$ and $\boldsymbol{v}\in V_{0}^{{\rm curl\,}}$, introduce the bi-linear form
$b_{h}(\boldsymbol{\tau},\psi;\boldsymbol{v}):=\langle({\rm curl\,}\operatorname{div})_{w}\boldsymbol{\tau},\boldsymbol{v}\rangle+({\rm grad\,}\psi,\boldsymbol{v}).$
By (15), we have for $(\boldsymbol{\tau},\psi)\in\Sigma_{k-1,h}^{\rm tn}\times\mathring{\mathbb{V}}_{\ell+1,h}^{{\rm grad\,}}$ and $\boldsymbol{v}\in\mathring{\mathbb{V}}_{(k,\ell),h}^{{\rm curl\,}}$ that
$b_{h}(\boldsymbol{\tau},\psi;\boldsymbol{v})=(({\rm curl\,}\operatorname{div})_{h}\boldsymbol{\tau},\boldsymbol{v})+({\rm grad\,}\psi,\boldsymbol{v}).$
Then the distributional mixed finite element method is to find $\boldsymbol{\sigma}_{h}\in{\Sigma}^{\rm tn}_{h}$, $\phi_{h}\in\mathring{\mathbb{V}}_{h}^{{\rm grad\,}}$ and $\boldsymbol{u}_{h}\in\mathring{\mathbb{V}}_{h}^{{\rm curl\,}}$ such that
(39a) $(\boldsymbol{\sigma}_{h},\boldsymbol{\tau}_{h})+b_{h}(\boldsymbol{\tau}_{h},\psi_{h};\boldsymbol{u}_{h})$ $=0,$ $\boldsymbol{\tau}_{h}\in{\Sigma}^{\rm tn}_{h},\psi_{h}\in\mathring{\mathbb{V}}_{h}^{{\rm grad\,}},$
(39b) $b_{h}(\boldsymbol{\sigma}_{h},\phi_{h};\boldsymbol{v}_{h})$ $=-\langle\boldsymbol{f},\boldsymbol{v}_{h}\rangle,$ $\boldsymbol{v}_{h}\in\mathring{\mathbb{V}}_{h}^{{\rm curl\,}}.$
We will derive two discrete inf-sup conditions for the linear form $b_{h}(\cdot,\cdot;\cdot)$. To this end, introduce some mesh dependent norms. For $\boldsymbol{\tau}\in\Sigma^{\rm tn}_{h}$, equip squared norm
$\|\boldsymbol{\tau}\|_{H^{-1}(({\rm curl\,}\operatorname{div})_{h})}^{2}$ $:=\|\boldsymbol{\tau}\|^{2}+\|({\rm curl\,}\operatorname{div})_{h}\boldsymbol{\tau}\|_{H_{h}^{-1}(\operatorname{div})}^{2},$
where $\|\boldsymbol{v}\|_{H_{h}^{-1}(\operatorname{div})}:=\sup_{\boldsymbol{w}_{h}\in\mathring{\mathbb{V}}_{h}^{{\rm curl\,}}}\frac{(\boldsymbol{v},\boldsymbol{w}_{h})}{\|\boldsymbol{w}_{h}\|_{H({\rm curl\,})}}$. The continuity of the bilinear form
$b_{h}(\boldsymbol{\tau},\psi;\boldsymbol{v})\leq(\|\boldsymbol{\tau}\|_{H^{-1}(({\rm curl\,}\operatorname{div})_{h})}+|\psi|_{1})\|\boldsymbol{v}\|_{H({\rm curl\,})},$
for all $\boldsymbol{\tau}\in{\Sigma}^{\rm tn},\psi\in\mathring{\mathbb{V}}_{h}^{{\rm grad\,}},\boldsymbol{v}\in\mathring{\mathbb{V}}_{h}^{{\rm curl\,}}$, is straight forward by the definition of these norms.
Lemma 4.4.
For $\boldsymbol{v}_{h}\in\mathring{\mathbb{V}}_{h}^{{\rm curl\,}}$, it holds
(40) $\|\boldsymbol{v}_{h}\|_{H({\rm curl\,})}\lesssim\sup_{\boldsymbol{\tau}_{h}\in{\Sigma}^{\rm tn}_{h},\psi_{h}\in\mathring{\mathbb{V}}_{h}^{{\rm grad\,}}}\frac{b_{h}(\boldsymbol{\tau}_{h},\psi_{h};\boldsymbol{v}_{h})}{\|\boldsymbol{\tau}_{h}\|_{H^{-1}(({\rm curl\,}\operatorname{div})_{h})}+|\psi_{h}|_{1}}.$
Proof.
By Helmholtz decomposition (29), given a $\boldsymbol{v}_{h}\in\mathring{\mathbb{V}}_{h}^{{\rm curl\,}}$, there exists $\boldsymbol{u}_{h}\in K_{h}^{c}$, $\tilde{\boldsymbol{v}}_{h}=({\rm curl\,}{\rm curl\,})_{h}\boldsymbol{u}_{h}$ and $\psi_{h}\in\mathring{\mathbb{V}}_{\ell+1,h}^{{\rm grad\,}}$ s.t.
(41) $\boldsymbol{v}_{h}=({\rm curl\,}{\rm curl\,})_{h}\boldsymbol{u}_{h}\oplus^{L_{2}}{\rm grad\,}\psi_{h}=\tilde{\boldsymbol{v}}_{h}\oplus^{L_{2}}{\rm grad\,}\psi_{h}.$
Then
${\rm curl\,}\boldsymbol{v}_{h}={\rm curl\,}\tilde{\boldsymbol{v}}_{h},\quad\|\boldsymbol{v}_{h}\|^{2}=\|({\rm curl\,}{\rm curl\,})_{h}\boldsymbol{u}_{h}\|^{2}+|\psi_{h}|_{1}^{2}.$
Set $\boldsymbol{\tau}_{h}=-I_{h}^{\rm tn}\operatorname{mskw}(\boldsymbol{u}_{h}+\tilde{\boldsymbol{v}}_{h})\in\Sigma_{k-1,h}^{\rm tn}$. By (25),
$(({\rm curl\,}\operatorname{div})_{h}\boldsymbol{\tau}_{h},\boldsymbol{v}_{h})=(({\rm curl\,}{\rm curl\,})_{h}(\boldsymbol{u}_{h}+\tilde{\boldsymbol{v}}_{h}),\boldsymbol{v}_{h})=\|({\rm curl\,}{\rm curl\,})_{h}\boldsymbol{u}_{h}\|^{2}+\|{\rm curl\,}\boldsymbol{v}_{h}\|^{2}.$
Consequently $b_{h}(\boldsymbol{\tau}_{h},\psi_{h};\boldsymbol{v}_{h})=\|\boldsymbol{v}_{h}\|^{2}+\|{\rm curl\,}\boldsymbol{v}_{h}\|^{2}$.
It remains to control the norms. As the decomposition (41) is $L^{2}$-orthogonal, $|\psi_{h}|_{1}\leq\|\boldsymbol{v}_{h}\|$ and $\|({\rm curl\,}{\rm curl\,})_{h}\boldsymbol{u}_{h}\|\leq\|\boldsymbol{v}_{h}\|$. We control the negative norm by
$\|({\rm curl\,}\operatorname{div})_{h}\boldsymbol{\tau}_{h}\|_{H_{h}^{-1}(\operatorname{div})}$ $=\sup_{\boldsymbol{w}_{h}\in\mathring{\mathbb{V}}_{h}^{{\rm curl\,}}}\frac{({\rm curl\,}\boldsymbol{u}_{h}+{\rm curl\,}\tilde{\boldsymbol{v}}_{h},{\rm curl\,}\boldsymbol{w}_{h})}{\|\boldsymbol{w}_{h}\|_{H({\rm curl\,})}}$
$\leq\|{\rm curl\,}\boldsymbol{u}_{h}\|+\|{\rm curl\,}\boldsymbol{v}_{h}\|$
$\lesssim\|({\rm curl\,}{\rm curl\,})_{h}\boldsymbol{u}_{h}\|+\|{\rm curl\,}\boldsymbol{v}_{h}\|\leq\|\boldsymbol{v}_{h}\|_{H({\rm curl\,})},$
where we have used the discrete Poincaré inequality (22).
By the scaling argument, the inverse inequality and the Poincaré inequality (22),
$\|\boldsymbol{\tau}_{h}\|\lesssim\|\boldsymbol{u}_{h}\|+\|\tilde{\boldsymbol{v}}_{h}\|\lesssim\|({\rm curl\,}{\rm curl\,})_{h}\boldsymbol{u}_{h}\|+\|{\rm curl\,}\boldsymbol{v}_{h}\|\leq 2\|\boldsymbol{v}_{h}\|_{H({\rm curl\,})},$
as required. ∎
Introduce
$\|\boldsymbol{v}_{h}\|_{H(({\rm grad\,}{\rm curl\,})_{h})}^{2}:=\|\boldsymbol{v}_{h}\|^{2}+\|({\rm grad\,}{\rm curl\,})_{h}\boldsymbol{v}_{h}\|^{2}.$
Again the continuity of the bilinear form in these norms
$b_{h}(\boldsymbol{\tau},\psi;\boldsymbol{v})\lesssim(\|\boldsymbol{\tau}\|+|\psi|_{1})\|\boldsymbol{v}\|_{H(({\rm grad\,}{\rm curl\,})_{h})},\boldsymbol{\tau}\in{\Sigma}^{\rm tn}_{h},\psi\in\mathring{\mathbb{V}}_{h}^{{\rm grad\,}},\boldsymbol{v}\in\mathring{\mathbb{V}}_{h}^{{\rm curl\,}}$
is straight forward by the definition of these norms.
Lemma 4.5.
For $\boldsymbol{v}_{h}\in\mathring{\mathbb{V}}_{h}^{{\rm curl\,}}$, it holds
(42) $\|\boldsymbol{v}_{h}\|_{H(({\rm grad\,}{\rm curl\,})_{h})}\lesssim\sup_{\boldsymbol{\tau}_{h}\in{\Sigma}^{\rm tn}_{h},\psi_{h}\in\mathring{\mathbb{V}}_{h}^{{\rm grad\,}}}\frac{b_{h}(\boldsymbol{\tau}_{h},\psi_{h};\boldsymbol{v}_{h})}{\|\boldsymbol{\tau}_{h}\|+|\psi_{h}|_{1}}.$
Proof.
We still use the Helmholtz decomposition (41) but choose
$\boldsymbol{\tau}_{h}=-({\rm grad\,}{\rm curl\,})_{h}\boldsymbol{v}_{h}=-({\rm grad\,}{\rm curl\,})_{h}\tilde{\boldsymbol{v}}_{h}.$
Then
$b_{h}(\boldsymbol{\tau}_{h},\psi_{h};\boldsymbol{v}_{h})=\|({\rm grad\,}{\rm curl\,})_{h}\boldsymbol{v}_{h}\|^{2}+|\psi_{h}|_{1}^{2}.$
We end the proof by the estimates $\|\boldsymbol{\tau}_{h}\|+|\psi_{h}|_{1}\lesssim\|\boldsymbol{v}_{h}\|_{H(({\rm grad\,}{\rm curl\,})_{h})}$ and $\|\boldsymbol{v}_{h}\|\lesssim\|({\rm grad\,}{\rm curl\,})_{h}\boldsymbol{v}_{h}\|+|\psi_{h}|_{1}$, in which we use the Poincaré inequality (30) for $\tilde{\boldsymbol{v}}_{h}$. ∎
There are other variants of mesh-dependent norms. For $\boldsymbol{\tau}\in\Sigma^{\rm tn}$, equip a mesh-dependent squared norm
$\|\boldsymbol{\tau}\|_{0,h}^{2}:=\|\boldsymbol{\tau}\|^{2}+\sum_{F\in\mathcal{F}_{h}}h_{F}\|\boldsymbol{n}\times\boldsymbol{\tau}\boldsymbol{n}\|_{F}^{2}.$
By the inverse trace inequality, clearly we have $\|\boldsymbol{\tau}_{h}\|_{0,h}\eqsim\|\boldsymbol{\tau}_{h}\|$ for $\boldsymbol{\tau}_{h}\in\Sigma^{\rm tn}_{h}$. For piecewise smooth vector-valued function $\boldsymbol{v}$, equip a mesh-dependent squared norm
$|\boldsymbol{v}|_{1,h}^{2}:=\sum_{T\in\mathcal{T}_{h}}\|{\rm grad\,}\boldsymbol{v}\|_{T}^{2}+\sum_{F\in\mathcal{F}_{h}}h_{F}^{-1}\|\llbracket\boldsymbol{v}\rrbracket\|_{F}^{2}.$
Then
$b_{h}(\boldsymbol{\tau},\psi;\boldsymbol{v})\lesssim\|\boldsymbol{\tau}\|_{0,h}|{\rm curl\,}\boldsymbol{v}|_{1,h}+|\psi|_{1}\|\boldsymbol{v}\|,\quad\boldsymbol{\tau}\in{\Sigma}^{\rm tn},\psi\in H_{0}^{1}(\Omega),\boldsymbol{v}\in V_{0}^{{\rm curl\,}}.$
One can prove the norm equivalence
(43) $\|({\rm grad\,}{\rm curl\,})_{h}\boldsymbol{v}_{h}\|\eqsim|{\rm curl\,}\boldsymbol{v}_{h}|_{1,h},\quad\boldsymbol{v}_{h}\in\mathring{\mathbb{V}}_{h}^{{\rm curl\,}}$
and thus obtain the discrete inf-sup condition from (42)
$\|{\rm curl\,}\boldsymbol{v}_{h}\|_{1,h}\lesssim\sup_{\boldsymbol{\tau}_{h}\in{\Sigma}^{\rm tn}_{h},\psi\in\mathring{\mathbb{V}}_{h}^{{\rm grad\,}}}\frac{b_{h}(\boldsymbol{\tau}_{h},\psi_{h};\boldsymbol{v}_{h})}{\|\boldsymbol{\tau}_{h}\|_{0,h}+|\psi_{h}|_{1}},$
where $\|{\rm curl\,}\boldsymbol{v}_{h}\|_{1,h}^{2}:=\|{\rm curl\,}\boldsymbol{v}_{h}\|^{2}+|{\rm curl\,}\boldsymbol{v}_{h}|_{1,h}^{2}$.
The discrete coercivity on the null space is similar to Lemma 4.2.
Lemma 4.6.
For $\boldsymbol{\tau}_{h}\in{\Sigma}^{\rm tn}_{h}$ and $\psi_{h}\in\mathring{\mathbb{V}}_{h}^{{\rm grad\,}}$ satisfying
$b_{h}(\boldsymbol{\tau}_{h},\psi_{h};\boldsymbol{v}_{h})=0,\quad\boldsymbol{v}_{h}\in\mathring{\mathbb{V}}_{h}^{{\rm curl\,}},$
it holds
(44) $\|\boldsymbol{\tau}_{h}\|_{H^{-1}(({\rm curl\,}\operatorname{div})_{h})}^{2}+|\psi_{h}|_{1}^{2}=\|\boldsymbol{\tau}_{h}\|^{2}.$
Applying the Babuška-Brezzi theory [5], from the discrete inf-sup conditions (40) and (42), and the discrete coercivity (44), we achieve the well-posedness of the mixed finite element method (39a)-(39b).
Theorem 4.7.
 The distributional mixed finite element method (39a)-(39b) for the quad-curl problem is well-posed. We have the discrete stability results
$\|{\boldsymbol{\sigma}}_{h}\|_{H^{-1}(({\rm curl\,}\operatorname{div})_{h})}+|{\phi}_{h}|_{1}+\|{\boldsymbol{u}}_{h}\|_{H({\rm curl\,})}$
(45) $\quad\lesssim\sup_{\boldsymbol{\tau}_{h}\in{\Sigma}^{\rm tn}_{h},\psi_{h}\in\mathring{\mathbb{V}}_{h}^{{\rm grad\,}},\boldsymbol{v}_{h}\in\mathring{\mathbb{V}}_{h}^{{\rm curl\,}}}\frac{A_{h}({\boldsymbol{\sigma}}_{h},{\phi}_{h},{\boldsymbol{u}}_{h};\boldsymbol{\tau}_{h},\psi_{h},\boldsymbol{v}_{h})}{\|\boldsymbol{\tau}_{h}\|_{H^{-1}(({\rm curl\,}\operatorname{div})_{h})}+|\psi_{h}|_{1}+\|\boldsymbol{v}_{h}\|_{H({\rm curl\,})}},$
$\|{\boldsymbol{\sigma}}_{h}\|+|{\phi}_{h}|_{1}+\|{\boldsymbol{u}}_{h}\|_{H(({\rm grad\,}{\rm curl\,})_{h})}$
(46) $\quad\lesssim\sup_{\boldsymbol{\tau}_{h}\in{\Sigma}^{\rm tn}_{h},\psi_{h}\in\mathring{\mathbb{V}}_{h}^{{\rm grad\,}},\boldsymbol{v}_{h}\in\mathring{\mathbb{V}}_{h}^{{\rm curl\,}}}\frac{A_{h}({\boldsymbol{\sigma}}_{h},{\phi}_{h},{\boldsymbol{u}}_{h};\boldsymbol{\tau}_{h},\psi_{h},\boldsymbol{v}_{h})}{\|\boldsymbol{\tau}_{h}\|+|\psi_{h}|_{1}+\|\boldsymbol{v}_{h}\|_{H(({\rm grad\,}{\rm curl\,})_{h})}},$
for any ${\boldsymbol{\sigma}}_{h}\in{\Sigma}^{\rm tn}_{h}$, ${\phi}_{h}\in\mathring{\mathbb{V}}_{h}^{{\rm grad\,}}$ and ${\boldsymbol{u}}_{h}\in\mathring{\mathbb{V}}_{h}^{{\rm curl\,}}$, where
$A_{h}({\boldsymbol{\sigma}}_{h},{\phi}_{h},{\boldsymbol{u}}_{h};\boldsymbol{\tau}_{h},\psi_{h},\boldsymbol{v}_{h}):=({\boldsymbol{\sigma}}_{h},\boldsymbol{\tau}_{h})+b_{h}(\boldsymbol{\tau}_{h},\psi_{h};{\boldsymbol{u}}_{h})+b_{h}({\boldsymbol{\sigma}}_{h},{\phi}_{h};\boldsymbol{v}_{h}).$
By choosing $\boldsymbol{v}_{h}={\rm grad\,}\phi_{h}$ in (39b), we get $\phi_{h}=0$ from $\operatorname{div}\boldsymbol{f}=0$.
4.3. Error analysis
Lemma 4.8.
Let $(\boldsymbol{\sigma},0,\boldsymbol{u})$ and $(\boldsymbol{\sigma}_{h},0,\boldsymbol{u}_{h})$ be the solution of the mixed formulation (34a)-(34b) and the mixed finite element method (39a)-(39b) respectively. Assume $\boldsymbol{\sigma}\in\Sigma^{\rm tn}$, and $\boldsymbol{u},{\rm curl\,}\boldsymbol{u}\in H^{1}(\Omega;\mathbb{R}^{3})$. Then
$\quad\;A_{h}(I_{h}^{\rm tn}\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h},0,I_{h}^{{\rm curl\,}}\boldsymbol{u}-\boldsymbol{u}_{h};\boldsymbol{\tau}_{h},\psi_{h},\boldsymbol{v}_{h})$
(47) $=(I_{h}^{\rm tn}\boldsymbol{\sigma}-\boldsymbol{\sigma},\boldsymbol{\tau}_{h})+(I_{h}^{{\rm curl\,}}\boldsymbol{u}-\boldsymbol{u},{\rm grad\,}\psi_{h})$
holds for any $\boldsymbol{\tau}_{h}\in{\Sigma}^{\rm tn}_{h}$, $\psi_{h}\in\mathring{\mathbb{V}}_{h}^{{\rm grad\,}}$ and $\boldsymbol{v}_{h}\in\mathring{\mathbb{V}}_{h}^{{\rm curl\,}}$.
Proof.
Subtract (39a)-(39b) from (34a)-(34b) and use (23) and (24) to get error equations
$(\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h},\boldsymbol{\tau}_{h})+b_{h}(\boldsymbol{\tau}_{h},\psi_{h};I_{h}^{{\rm curl\,}}\boldsymbol{u}-\boldsymbol{u}_{h})$ $=(I_{h}^{{\rm curl\,}}\boldsymbol{u}-\boldsymbol{u},{\rm grad\,}\psi_{h}),$
$b_{h}(I_{h}^{\rm tn}\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h},0;\boldsymbol{v}_{h})$ $=0.$
Then subtract $(\boldsymbol{\sigma}-I_{h}^{\rm tn}\boldsymbol{\sigma},\boldsymbol{\tau}_{h})$ to get (47). ∎
Theorem 4.9.
Let $(\boldsymbol{\sigma},0,\boldsymbol{u})$ and $(\boldsymbol{\sigma}_{h},0,\boldsymbol{u}_{h})$ be the solution of the mixed formulation (34a)-(34b) and the mixed finite element method (39a)-(39b) respectively. Assume $\boldsymbol{\sigma}\in H^{k}(\Omega;\mathbb{T})$ and $\boldsymbol{u},{\rm curl\,}\boldsymbol{u}\in H^{k}(\Omega;\mathbb{R}^{3})$. Then
(48) $\|\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h}\|+\|I_{h}^{{\rm curl\,}}\boldsymbol{u}-\boldsymbol{u}_{h}\|_{H(({\rm grad\,}{\rm curl\,})_{h})}$ $\lesssim h^{k}(|\boldsymbol{\sigma}|_{k}+|\boldsymbol{u}|_{k}),$
(49) $\|\boldsymbol{u}-\boldsymbol{u}_{h}\|_{H({\rm curl\,})}+h|{\rm curl\,}(\boldsymbol{u}-\boldsymbol{u}_{h})|_{1,h}$ $\lesssim h^{k}(|\boldsymbol{\sigma}|_{k}+|\boldsymbol{u}|_{k}+|{\rm curl\,}\boldsymbol{u}|_{k}).$
Proof.
It follows from the stability results (45)-(46) and (47) that
$\quad\|I_{h}^{\rm tn}\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h}\|+\|I_{h}^{{\rm curl\,}}\boldsymbol{u}-\boldsymbol{u}_{h}\|_{H({\rm curl\,})}+\|I_{h}^{{\rm curl\,}}\boldsymbol{u}-\boldsymbol{u}_{h}\|_{H(({\rm grad\,}{\rm curl\,})_{h})}$
$\lesssim\|\boldsymbol{\sigma}-I_{h}^{\rm tn}\boldsymbol{\sigma}\|+\|\boldsymbol{u}-I_{h}^{{\rm curl\,}}\boldsymbol{u}\|.$
Hence (48)-(49) follow from the triangle inequality, the norm equivalence (43) and interpolation error estimates. ∎
Using the usual duality argument, we can additionally derive the superconvergence of $\|{\rm curl\,}(I_{h}^{{\rm curl\,}}\boldsymbol{u}-\boldsymbol{u}_{h})\|_{0}$. Let $\widetilde{\boldsymbol{u}}$ be the solution of the auxiliary problem:
(50) {$-{\rm curl\,}\Delta{\rm curl\,}\widetilde{\boldsymbol{u}}&=-{\rm curl\,}^{2}(I_{h}^{{\rm curl\,}}\boldsymbol{u}-\boldsymbol{u}_{h})&&\text{ in }\Omega,\\ \operatorname{div}\widetilde{\boldsymbol{u}}&=0&&\text{ in }\Omega,\\ \widetilde{\boldsymbol{u}}\times\boldsymbol{n}=\operatorname{curl}\widetilde{\boldsymbol{u}}\times\boldsymbol{n}&=0&&\text{ on }\partial\Omega.\end{aligned}\right.$
We assume that $\widetilde{\boldsymbol{u}}\in H^{1}(\Omega;\mathbb{R}^{3})\cap H_{0}({\rm curl\,},\Omega)$ with the bound
(51) $\|\widetilde{\boldsymbol{\sigma}}\|_{1}+\|\widetilde{\boldsymbol{u}}\|_{1}\lesssim\|{\rm curl\,}(I_{h}^{{\rm curl\,}}\boldsymbol{u}-\boldsymbol{u}_{h})\|,$
where $\widetilde{\boldsymbol{\sigma}}:={\rm grad\,}{\rm curl\,}\widetilde{\boldsymbol{u}}$. When $\Omega$ is convex, regularity (51) has been obtained in [28, Lemma A.1].
Theorem 4.10.
Let $(\boldsymbol{\sigma},0,\boldsymbol{u})$ and $(\boldsymbol{\sigma}_{h},0,\boldsymbol{u}_{h})$ be the solution of the mixed formulation (34a)-(34b) and the mixed finite element method (39a)-(39b) respectively. Assume the regularity condition (51) holds, $\boldsymbol{f}\in L^{2}(\Omega;\mathbb{R}^{3})$, $\boldsymbol{\sigma}\in H^{k}(\Omega;\mathbb{T})$ and $\boldsymbol{u},{\rm curl\,}\boldsymbol{u}\in H^{k}(\Omega;\mathbb{R}^{3})$. Then
(52) $\|{\rm curl\,}(I_{h}^{{\rm curl\,}}\boldsymbol{u}-\boldsymbol{u}_{h})\|\lesssim h^{k+1}(|\boldsymbol{\sigma}|_{k}+|\boldsymbol{u}|_{k}+\delta_{k1}\|\boldsymbol{f}\|),$
where $\delta_{11}=1$ and $\delta_{k1}=0$ for $k\neq 1$.
Proof.
By the first equation of (50), we have
$\|{\rm curl\,}(I_{h}^{{\rm curl\,}}\boldsymbol{u}-\boldsymbol{u}_{h})\|^{2}=(\operatorname{div}\widetilde{\boldsymbol{\sigma}},{\rm curl\,}(I_{h}^{{\rm curl\,}}\boldsymbol{u}-\boldsymbol{u}_{h}))=b_{h}(\widetilde{\boldsymbol{\sigma}},0;I_{h}^{{\rm curl\,}}\boldsymbol{u}-\boldsymbol{u}_{h}).$
Then we get from (24) and (23) that
$\|{\rm curl\,}(I_{h}^{{\rm curl\,}}\boldsymbol{u}-\boldsymbol{u}_{h})\|^{2}=b_{h}(I_{h}^{\rm tn}\widetilde{\boldsymbol{\sigma}},0;I_{h}^{{\rm curl\,}}\boldsymbol{u}-\boldsymbol{u}_{h})=b_{h}(I_{h}^{\rm tn}\widetilde{\boldsymbol{\sigma}},0;\boldsymbol{u}-\boldsymbol{u}_{h}),$
which together with (34a) and (39a) gives
$\|{\rm curl\,}(I_{h}^{{\rm curl\,}}\boldsymbol{u}-\boldsymbol{u}_{h})\|^{2}=-(\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h},I_{h}^{\rm tn}\widetilde{\boldsymbol{\sigma}}).$
On the other side, it follows from (34b), (39b) and (24) that
$-(\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h},\widetilde{\boldsymbol{\sigma}})$ $=-(\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h},{\rm grad\,}{\rm curl\,}\widetilde{\boldsymbol{u}})=b_{h}(\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h},0;\widetilde{\boldsymbol{u}})$
$=b_{h}(\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h},0;\widetilde{\boldsymbol{u}}-I_{h}^{{\rm curl\,}}\widetilde{\boldsymbol{u}})=b_{h}(\boldsymbol{\sigma},0;\widetilde{\boldsymbol{u}}-I_{h}^{{\rm curl\,}}\widetilde{\boldsymbol{u}}).$
Combining the last two equations and the estimate of $I_{h}^{\rm tn}$ yields
$\|{\rm curl\,}(I_{h}^{{\rm curl\,}}\boldsymbol{u}-\boldsymbol{u}_{h})\|^{2}$ $=(\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h},\widetilde{\boldsymbol{\sigma}}-I_{h}^{\rm tn}\widetilde{\boldsymbol{\sigma}})+b_{h}(\boldsymbol{\sigma},0;\widetilde{\boldsymbol{u}}-I_{h}^{{\rm curl\,}}\widetilde{\boldsymbol{u}})$
(53) $\lesssim h\|\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h}\|\|\widetilde{\boldsymbol{\sigma}}\|_{1}+b_{h}(\boldsymbol{\sigma},0;\widetilde{\boldsymbol{u}}-I_{h}^{{\rm curl\,}}\widetilde{\boldsymbol{u}}).$
When $k\geq 2$, applying (24) again, we get from (53) that
$\|{\rm curl\,}(I_{h}^{{\rm curl\,}}\boldsymbol{u}-\boldsymbol{u}_{h})\|^{2}$ $\lesssim h\|\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h}\|\|\widetilde{\boldsymbol{\sigma}}\|_{1}+b_{h}(\boldsymbol{\sigma}-I_{h}^{\rm tn}\boldsymbol{\sigma},0;\widetilde{\boldsymbol{u}}-I_{h}^{{\rm curl\,}}\widetilde{\boldsymbol{u}})$
$\leq h\|\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h}\|\|\widetilde{\boldsymbol{\sigma}}\|_{1}$
$\quad+\|\boldsymbol{\sigma}-I_{h}^{\rm tn}\boldsymbol{\sigma}\|_{0,h}|{\rm curl\,}(\widetilde{\boldsymbol{u}}-I_{h}^{{\rm curl\,}}\widetilde{\boldsymbol{u}})|_{1,h}$
$\lesssim h(\|\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h}\|+\|\boldsymbol{\sigma}-I_{h}^{\rm tn}\boldsymbol{\sigma}\|_{0,h})\|\widetilde{\boldsymbol{\sigma}}\|_{1}.$
Therefore (52) follows from (48), interpolation error estimate of $I_{h}^{\rm tn}$ and (51).
Next consider case $k=1$. By the de Rham complex [18], there exists $\boldsymbol{w}\in H_{0}^{1}(\Omega;\mathbb{R}^{3})$ such that
${\rm curl\,}\boldsymbol{w}={\rm curl\,}(\widetilde{\boldsymbol{u}}-I_{h}^{{\rm curl\,}}\widetilde{\boldsymbol{u}}),\quad\|\boldsymbol{w}\|_{1}\lesssim\|{\rm curl\,}\boldsymbol{w}\|\lesssim h|{\rm curl\,}\widetilde{\boldsymbol{u}}|_{1}.$
Then apply the commutative property of $I_{h}^{{\rm curl\,}}$ to get ${\rm curl\,}(I_{h}^{{\rm curl\,}}\boldsymbol{w})=0$. Employing the second equation of problem (33) and the estimate of $I_{h}^{{\rm curl\,}}$, we acquire
$b_{h}(\boldsymbol{\sigma},0;\widetilde{\boldsymbol{u}}-I_{h}^{{\rm curl\,}}\widetilde{\boldsymbol{u}})$ $=(\operatorname{div}\boldsymbol{\sigma},{\rm curl\,}(\widetilde{\boldsymbol{u}}-I_{h}^{{\rm curl\,}}\widetilde{\boldsymbol{u}}))=(\operatorname{div}\boldsymbol{\sigma},{\rm curl\,}({\boldsymbol{w}}-I_{h}^{{\rm curl\,}}{\boldsymbol{w}}))$
$=({\rm curl\,}\operatorname{div}\boldsymbol{\sigma},{\boldsymbol{w}}-I_{h}^{{\rm curl\,}}{\boldsymbol{w}})=-(\boldsymbol{f},{\boldsymbol{w}}-I_{h}^{{\rm curl\,}}{\boldsymbol{w}})$
$\lesssim h\|\boldsymbol{f}\|\|\boldsymbol{w}\|_{1}\lesssim h^{2}\|\boldsymbol{f}\||{\rm curl\,}\widetilde{\boldsymbol{u}}|_{1}.$
Finally, (52) follows from (53), (48) and (51). ∎
We perform numerical experiments to support the theoretical results of the distributional mixed method (39a)-(39b). Let $\Omega=(0,1)^{3}$. Choose the function $\boldsymbol{f}$ in (32) such that the exact solution of (32) is
$\boldsymbol{u}={\rm curl\,}\begin{pmatrix}\sin^{3}(\pi x)\sin^{3}(\pi y)\sin^{3}(\pi z)\\ \sin^{3}(\pi x)\sin^{3}(\pi y)\sin^{3}(\pi z)\\ 0\end{pmatrix},$
and let $\boldsymbol{\sigma}:={\rm grad\,}{\rm curl\,}\boldsymbol{u}$. To break the symmetry, the initial unstructured mesh of $\Omega$ is shown in Fig. 4, which is a perturbation of the uniform mesh with $h=1/2$. Then we take uniform refinement of this initial unstructured triangulation.
Figure 4. An initial perturbed mesh of the uniform mesh with $h=1/2$.
We will implement the hybridized version of (39a)-(39b); see (60a)-(60b). Therefore the number of DoFs is $\dim\mathring{\mathbb{V}}^{{\rm curl\,}}_{(k,\ell),h}+\dim\Lambda_{k-1,h}$, where $\Lambda_{k-1,h}$ is the space of Lagrange multiplier. Numerical results of $\|\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h}\|$, $\|{\boldsymbol{u}}-\boldsymbol{u}_{h}\|$, $\|{\rm curl\,}({\boldsymbol{u}}-\boldsymbol{u}_{h})\|$ and $\|{\rm grad\,}_{\mathcal{T}_{h}}{\rm curl\,}(\boldsymbol{u}-\boldsymbol{u}_{h})\|$ of the distributional mixed method (39a)-(39b) with $k=1$ and $\ell=0,1$ are shown in Table 1. It is observed that $\|\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h}\|=\mathcal{O}(h)$, and $\|{\rm curl\,}({\boldsymbol{u}}-\boldsymbol{u}_{h})\|=\mathcal{O}(h)$ numerically, which coincide with the theoretical error estimates in (48) and (49). The error $\|{\boldsymbol{u}}-\boldsymbol{u}_{h}\|=\mathcal{O}(h^{\ell+1})$ is also observed but not included in (48) and (49).
Table 1. Errors for the distributional mixed finite element method (39a)-(39b) with $k=1$.
$h$ $2^{-2}$ $2^{-3}$ $2^{-4}$ $2^{-5}$
# DoFs for $\ell=0$ 2,332 17,240 132,400 1,037,408
# DoFs for $\ell=1$ 2,936 21,424 163,424 1,276,096
$\|\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h}\|$ 1.51E+02 8.62E+01 4.48E+01 2.26E+01
order -- 0.81 0.95 0.99
$\|{\rm curl\,}(\boldsymbol{u}-\boldsymbol{u}_{h})\|$ 1.48E+01 5.60E+00 2.39E+00 1.13E+00
order -- 1.40 1.23 1.08
$\|{\rm grad\,}_{\mathcal{T}_{h}}{\rm curl\,}(\boldsymbol{u}-\boldsymbol{u}_{h})\|$ 1.46E+02 1.46E+02 1.46E+02 1.46E+02
order -- 0 0 0
$\|\boldsymbol{u}-\boldsymbol{u}_{h}\|$ for $\ell=0$ 1.49E+00 5.23E-01 2.19E-01 1.03E-01
order -- 1.51 1.25 1.09
$\|\boldsymbol{u}-\boldsymbol{u}_{h}\|$ for $\ell=1$ 1.04E+00 2.82E-01 7.32E-02 1.85E-02
order -- 1.88 1.95 1.98
4.4. Post-processing
Post-processing can be also applied to improve the approximation. It follows from the standard procedure of the stable mixed methods and will be briefly reviewed below.
We will construct a new superconvergent approximation to $\boldsymbol{u}$ in virtue of the optimal result of $\boldsymbol{\sigma}$ in (48). Introduce discrete space
$\mathbb{V}_{h}^{\ast}$ :=$:=\{\boldsymbol{v}_{h}\in L^{2}(\Omega;\mathbb{R}^{3}):\boldsymbol{v}_{h}|_{T}\in\mathbb{P}_{k}(T;\mathbb{R}^{3})+\boldsymbol{x}\times\mathbb{P}_{k}(T;\mathbb{R}^{3})\textrm{ for }T\in\mathcal{T}_{h}\}.$
Let ̊$\mathring{\mathbb{P}}_{k+1}(T)$ be the subspace of $\mathbb{P}_{k+1}(T)$ with vanishing funcition values at vertices of $T$. For each $T\in\mathcal{T}_{h}$, define the new approximation $\boldsymbol{u}_{h}^{\ast}\in\mathbb{V}_{h}^{\ast}$ to $\boldsymbol{u}$ piecewisely as a solution of the following problem:
(54) $(\boldsymbol{u}_{h}^{\ast}\cdot\boldsymbol{t},q)_{e}$ $=(\boldsymbol{u}_{h}\cdot\boldsymbol{t},q)_{e}\quad\;\forall~{}q\in\mathbb{P}_{0}(e),e\in\mathcal{E}(T),$
(55) $({\rm grad\,}{\rm curl\,}\boldsymbol{u}_{h}^{\ast},\boldsymbol{q})_{T}$ $=(\boldsymbol{\sigma}_{h},\boldsymbol{q})_{T}\qquad\forall~{}\boldsymbol{q}\in{\rm grad\,}{\rm curl\,}\mathbb{P}_{k+1}(T;\mathbb{R}^{3}),$
(56) $(\boldsymbol{u}_{h}^{\ast},\boldsymbol{q})_{T}$ $=(\boldsymbol{u}_{h},\boldsymbol{q})_{T}\qquad\forall~{}\boldsymbol{q}\in{\rm grad\,}\mathring{\mathbb{P}}_{k+1}(T).$
It is easy to verify that the local problem (54)-(56) is well-posed.
Theorem 4.11.
Let $(\boldsymbol{\sigma},0,\boldsymbol{u})$ and $(\boldsymbol{\sigma}_{h},0,\boldsymbol{u}_{h})$ be the solution of the mixed formulation (34a)-(34b) and the mixed finite element method (39a)-(39b) respectively. Assume $\boldsymbol{\sigma}\in H^{k}(\Omega;\mathbb{T})$ and $\boldsymbol{u},{\rm curl\,}\boldsymbol{u}\in H^{k+1}(\Omega;\mathbb{R}^{3})$. Then
(57) $\|{\rm grad\,}_{\mathcal{T}_{h}}{\rm curl\,}_{\mathcal{T}_{h}}(\boldsymbol{u}-\boldsymbol{u}_{h}^{\ast})\|\lesssim h^{k}(|\boldsymbol{\sigma}|_{k}+|\boldsymbol{u}|_{k}+|{\rm curl\,}\boldsymbol{u}|_{k+1}),$
Further assume the regularity condition (51) holds and $\boldsymbol{f}\in L^{2}(\Omega;\mathbb{R}^{3})$. Then
(58) $\|{\rm curl\,}_{\mathcal{T}_{h}}(\boldsymbol{u}-\boldsymbol{u}_{h}^{\ast})\|\lesssim h^{k+1}(|\boldsymbol{\sigma}|_{k}+|\boldsymbol{u}|_{k}+|{\rm curl\,}\boldsymbol{u}|_{k+1}+\delta_{k1}\|\boldsymbol{f}\|).$
Proof.
Denote $\boldsymbol{w}=(I-I_{T,1}^{{\rm curl\,}})(I_{T,k+1}^{{\rm curl\,}}\boldsymbol{u}-\boldsymbol{u}_{h}^{\ast})\in\mathbb{P}_{k}(T;\mathbb{R}^{3})+\boldsymbol{x}\times\mathbb{P}_{k}(T;\mathbb{R}^{3})$. From the first equation of problem (33) and (55) with $\boldsymbol{q}={\rm grad\,}{\rm curl\,}\boldsymbol{w}$, it follows that
$|{\rm curl\,}\boldsymbol{w}|_{1,T}^{2}$ $=({\rm grad\,}{\rm curl\,}(I_{T,k+1}^{{\rm curl\,}}\boldsymbol{u}-\boldsymbol{u}_{h}^{\ast}),{\rm grad\,}{\rm curl\,}\boldsymbol{w})_{T}$
$=({\rm grad\,}{\rm curl\,}(I_{T,k+1}^{{\rm curl\,}}\boldsymbol{u}-\boldsymbol{u}),{\rm grad\,}{\rm curl\,}\boldsymbol{w})_{T}+(\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h},{\rm grad\,}{\rm curl\,}\boldsymbol{w})_{T},$
which means
(59) $|{\rm curl\,}\boldsymbol{w}|_{1,T}\leq|{\rm curl\,}(I_{T,k+1}^{{\rm curl\,}}\boldsymbol{u}-\boldsymbol{u})|_{1,T}+\|\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h}\|_{0,T}.$
Then
$|{\rm curl\,}(\boldsymbol{u}-\boldsymbol{u}_{h}^{\ast})|_{1,T}\leq 2|{\rm curl\,}(I_{T,k+1}^{{\rm curl\,}}\boldsymbol{u}-\boldsymbol{u})|_{1,T}+\|\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h}\|_{0,T}.$
Hence (57) follows from (48) and the error estimate of $I_{T,k+1}^{{\rm curl\,}}$.
Next we derive (58). It is easy to see that $I_{T,1}^{{\rm curl\,}}\boldsymbol{w}=0$. By the error estimate of $I_{T,1}^{{\rm curl\,}}$,
$\|{\rm curl\,}\boldsymbol{w}\|_{0,T}=\|{\rm curl\,}(\boldsymbol{w}-I_{T,1}^{{\rm curl\,}}\boldsymbol{w})\|_{0,T}\lesssim h_{K}|{\rm curl\,}\boldsymbol{w}|_{1,T},$
which together with (59) gives
$\|{\rm curl\,}\boldsymbol{w}\|_{0,T}\lesssim h_{K}|{\rm curl\,}\boldsymbol{w}|_{1,T}\lesssim h_{T}|{\rm curl\,}(I_{T,k+1}^{{\rm curl\,}}\boldsymbol{u}-\boldsymbol{u})|_{1,T}+h_{T}\|\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h}\|_{0,T}.$
On the other side, by (54), the commutative property of $I_{T,1}^{{\rm curl\,}}$ and the norm equivalence, we have
$\|{\rm curl\,}\big{(}I_{T,1}^{{\rm curl\,}}(I_{T,k+1}^{{\rm curl\,}}\boldsymbol{u}-\boldsymbol{u}_{h}^{\ast})\big{)}\|_{0,T}$ $=\|{\rm curl\,}\big{(}I_{T,1}^{{\rm curl\,}}(I_{h}^{{\rm curl\,}}\boldsymbol{u}-\boldsymbol{u}_{h}))\|_{0,T}$
$\lesssim\|{\rm curl\,}(I_{h}^{{\rm curl\,}}\boldsymbol{u}-\boldsymbol{u}_{h})\|_{0,T}.$
Then we get from the last two inequalities that
$\|{\rm curl\,}(\boldsymbol{u}-\boldsymbol{u}_{h}^{\ast})\|_{0,T}$ $\leq\|{\rm curl\,}(\boldsymbol{u}-I_{T,k+1}^{{\rm curl\,}}\boldsymbol{u})\|_{0,T}+\|{\rm curl\,}\boldsymbol{w}\|_{0,T}$
$\quad+\|{\rm curl\,}(I_{T,1}^{{\rm curl\,}}(I_{T,k+1}^{{\rm curl\,}}\boldsymbol{u}-\boldsymbol{u}_{h}^{\ast}))\|_{0,T}$
$\lesssim\|{\rm curl\,}(\boldsymbol{u}-I_{T,k+1}^{{\rm curl\,}}\boldsymbol{u})\|_{0,T}+h_{T}|{\rm curl\,}(I_{T,k+1}^{{\rm curl\,}}\boldsymbol{u}-\boldsymbol{u})|_{1,T}$
$\quad+h_{T}\|\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h}\|_{0,T}+\|{\rm curl\,}(I_{h}^{{\rm curl\,}}\boldsymbol{u}-\boldsymbol{u}_{h})\|_{0,T}.$
Finally, estimate (58) follows from the error estimate of $I_{T,k+1}^{{\rm curl\,}}$, (48) and (52). ∎
Table 2. Errors for the post-processing for $k=1$.
$h$ $2^{-2}$ $2^{-3}$ $2^{-4}$ $2^{-5}$
# DoFs for $\ell=0$ 2,332 17,240 132,400 1,037,408
# DoFs for $\ell=1$ 2,936 21,424 163,424 1,276,096
$\|{\rm curl\,}_{\mathcal{T}_{h}}(\boldsymbol{u}-\boldsymbol{u}_{h}^{\ast})\|$ 9.25E+00 2.71E+00 7.18E-01 1.83E-01
order -- 1.77 1.92 1.98
$\|{\rm grad\,}_{\mathcal{T}_{h}}{\rm curl\,}_{\mathcal{T}_{h}}(\boldsymbol{u}-\boldsymbol{u}_{h}^{\ast})\|$ 1.52E+02 8.62E+01 4.48E+01 2.26E+01
order -- 0.81 0.95 0.99
Numerical results of the post-processing $\boldsymbol{u}_{h}^{\ast}$ for $k=1$ and $\ell=0,1$ are listed in Table 2. We can observe that $\|{\rm curl\,}_{\mathcal{T}_{h}}(\boldsymbol{u}-\boldsymbol{u}_{h}^{\ast})\|=\mathcal{O}(h^{2})$ and $\|{\rm grad\,}_{\mathcal{T}_{h}}{\rm curl\,}_{\mathcal{T}_{h}}(\boldsymbol{u}-\boldsymbol{u}_{h}^{\ast})\|=\mathcal{O}(h)$ numerically, which are one order higher than $\|{\rm curl\,}({\boldsymbol{u}}-\boldsymbol{u}_{h})\|$ and $\|{\rm grad\,}_{\mathcal{T}_{h}}{\rm curl\,}({\boldsymbol{u}}-\boldsymbol{u}_{h})\|$ respectively. Indeed, $\|{\rm grad\,}_{\mathcal{T}_{h}}{\rm curl\,}({\boldsymbol{u}}-\boldsymbol{u}_{h})\|=\|\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h}\|$.
5. Hybridization of distributional mixed finite element method
In this section we will hybridize the distributional mixed finite element method (39a)-(39b) following the framework in [1]. We introduce a Lagrange multiplier for the tangential-normal continuity of $\boldsymbol{\sigma}$ which can be treat as an approximation of $(\boldsymbol{n}_{F}\times{\rm curl\,}\boldsymbol{u})|_{\mathcal{F}_{h}}$. As $\boldsymbol{\sigma}$ is discontinuous, it can be eliminated element-wise and the size of the resulting linear system is reduced, which is easier to solve than the saddle point system obtained by the mixed method (39a)-(39b). With the hybridized method, we can also establish the equivalence of the mixed method (39a)-(39b) to a weak Galerkin method without stabilization and nonconforming finite element methods in [28, 45].
To this end, introduce two finite element spaces
${\Sigma}_{k-1,h}^{-1}:=$$\{\boldsymbol{\tau}_{h}\in L^{2}(\Omega;\mathbb{T}):\boldsymbol{\tau}_{h}|_{T}\in\mathbb{P}_{k-1}(T;\mathbb{T})\text{ for each }T\in\mathcal{T}_{h}\},$
$\Lambda_{k-1,h}:=$$\{\boldsymbol{\mu}_{h}\in L^{2}(\mathcal{F}_{h};\mathbb{R}^{3}):\boldsymbol{\mu}_{h}|_{F}\in\mathbb{P}_{k-1}(F;\mathbb{R}^{3})\text{ and }\boldsymbol{\mu}_{h}\cdot\boldsymbol{n}|_{F}=0\text{ for each }F\in\mathring{\mathcal{F}}_{h},$
[…]
The space $\Lambda_{h}$ is introduced as the Lagrange multiplier to impose the tangential-normal continuity and is equipped with squared norm
$\left\|\boldsymbol{\mu}_{h}\right\|_{\alpha,h}^{2}:=\sum_{T\in\mathcal{T}_{h}}\sum_{F\in\mathcal{F}(T)}h_{F}^{-2\alpha}\left\|\boldsymbol{\mu}_{h}\right\|_{F}^{2},\quad\alpha=\pm 1/2.$
5.1. Hybridization
The hybridization of the mixed finite element method (39a)-(39b) is to find $(\boldsymbol{\sigma}_{h},\boldsymbol{u}_{h},\phi_{h},\boldsymbol{\lambda}_{h})\in{\Sigma}_{k-1,h}^{-1}\times\mathring{\mathbb{V}}^{{\rm curl\,}}_{(k,\ell),h}\times\mathring{\mathbb{V}}^{{\rm grad\,}}_{\ell+1,h}\times\Lambda_{k-1,h}$ such that
(60a) $(\boldsymbol{\sigma}_{h},\boldsymbol{\tau}_{h})+b_{h}(\boldsymbol{\tau}_{h},\psi_{h};\boldsymbol{u}_{h})\!+c_{h}(\boldsymbol{\tau}_{h},\boldsymbol{\lambda}_{h})$ $=0,\quad\;\;\;\;\;\;\;\boldsymbol{\tau}_{h}\in{\Sigma}_{k-1,h}^{-1},\psi_{h}\in\mathring{\mathbb{V}}^{{\rm grad\,}}_{\ell+1,h},$
(60b) $b_{h}(\boldsymbol{\sigma}_{h},\phi_{h};\boldsymbol{v}_{h})\!+c_{h}(\boldsymbol{\sigma}_{h},\boldsymbol{\mu}_{h})$ $=\!\!-\langle\boldsymbol{f},\boldsymbol{v}_{h}\rangle,\boldsymbol{v}_{h}\in\mathring{\mathbb{V}}^{{\rm curl\,}}_{(k,\ell),h},\boldsymbol{\mu}_{h}\in\Lambda_{k-1,h},$
where the bilinear form $c_{h}(\boldsymbol{\tau}_{h},\boldsymbol{\lambda}_{h}):=-\sum_{T\in\mathcal{T}_{h}}\left(\boldsymbol{n}\times\boldsymbol{\tau}_{h}\boldsymbol{n}_{F},\boldsymbol{\lambda}_{h}\right)_{\partial T}$ is introduced to impose the tangential-normal continuity.
Lemma 5.1.
There holds the following inf-sup condition
(61) $\begin{split}&\|\boldsymbol{v}_{h}\|_{H(({\rm grad\,}{\rm curl\,})_{h})}+\left\|\boldsymbol{n}_{F}\times{\rm curl\,}\boldsymbol{v}_{h}-\boldsymbol{\mu}_{h}\right\|_{1/2,h}\\ \lesssim&\sup_{\boldsymbol{\tau}_{h}\in\Sigma_{k-1,h}^{-1},\psi_{h}\in\mathring{\mathbb{V}}_{\ell+1,h}^{{\rm grad\,}}}\frac{b_{h}\left(\boldsymbol{\tau}_{h},\psi_{h};\boldsymbol{v}_{h}\right)+c_{h}(\boldsymbol{\tau}_{h},\boldsymbol{\mu}_{h})}{\left\|\boldsymbol{\tau}_{h}\right\|+|\psi_{h}|_{1}},\ \boldsymbol{v}_{h}\in\mathring{\mathbb{V}}_{(k,\ell),h}^{{\rm curl\,}},\boldsymbol{\mu}_{h}\in\Lambda_{k-1,h}.\end{split}$
Proof.
Let $\boldsymbol{\tau}_{h}\in\Sigma_{k-1,h}^{-1}$ be determined as follows: for $T\in\mathcal{T}_{h}$,
$(\boldsymbol{n}\times\boldsymbol{\tau}_{h}\boldsymbol{n}_{F})|_{F}$ $=\frac{1}{h_{F}}(\boldsymbol{n}_{F}\times{\rm curl\,}\boldsymbol{v}_{h}-\boldsymbol{\mu}_{h}),\qquad F\in\mathcal{F}(T),$
$(\boldsymbol{\tau}_{h},\boldsymbol{q})_{T}$ $=-({\rm grad\,}{\rm curl\,}\boldsymbol{v}_{h},\boldsymbol{q})_{T},\qquad\quad\;\;\boldsymbol{q}\in\mathbb{P}_{k-2}(T;\mathbb{T}).$
Clearly we have
$\|\boldsymbol{\tau}_{h}\|\lesssim\|{\rm grad\,}_{\mathcal{T}_{h}}{\rm curl\,}\boldsymbol{v}_{h}\|+\left\|\boldsymbol{n}_{F}\times{\rm curl\,}\boldsymbol{v}_{h}-\boldsymbol{\mu}_{h}\right\|_{1/2,h},$
$b_{h}\left(\boldsymbol{\tau}_{h},0;\boldsymbol{v}_{h}\right)+c_{h}(\boldsymbol{\tau}_{h},\boldsymbol{\mu}_{h})=\|{\rm grad\,}_{\mathcal{T}_{h}}{\rm curl\,}\boldsymbol{v}_{h}\|^{2}+\left\|\boldsymbol{n}_{F}\times{\rm curl\,}\boldsymbol{v}_{h}-\boldsymbol{\mu}_{h}\right\|_{1/2,h}^{2}.$
Then
(62) $\quad\|{\rm grad\,}_{\mathcal{T}_{h}}{\rm curl\,}\boldsymbol{v}_{h}\|+\left\|\boldsymbol{n}_{F}\times{\rm curl\,}\boldsymbol{v}_{h}-\boldsymbol{\mu}_{h}\right\|_{1/2,h}$
$\qquad\qquad\qquad\qquad\qquad\qquad\lesssim\sup_{\boldsymbol{\tau}_{h}\in\Sigma_{k-1,h}^{-1}}\frac{b_{h}\left(\boldsymbol{\tau}_{h},0;\boldsymbol{v}_{h}\right)+c_{h}(\boldsymbol{\tau}_{h},\boldsymbol{\mu}_{h})}{\left\|\boldsymbol{\tau}_{h}\right\|},$
which together with (43) indicates
$\quad\;\|({\rm grad\,}{\rm curl\,})_{h}\boldsymbol{v}_{h}\|+\left\|\boldsymbol{n}_{F}\times{\rm curl\,}\boldsymbol{v}_{h}-\boldsymbol{\mu}_{h}\right\|_{1/2,h}$
$\lesssim\sup_{\boldsymbol{\tau}_{h}\in\Sigma_{h}^{-1},\psi_{h}\in\mathring{\mathbb{V}}_{h}^{{\rm grad\,}}}\frac{b_{h}\left(\boldsymbol{\tau}_{h},\psi_{h};\boldsymbol{v}_{h}\right)+c_{h}(\boldsymbol{\tau}_{h},\boldsymbol{\mu}_{h})}{\left\|\boldsymbol{\tau}_{h}\right\|+|\psi_{h}|_{1}}.$
Therefore the inf-sup condition (LABEL:hybridizationinfsup) holds from (42). ∎
Theorem 5.2.
The hybridized mixed finite element method (60a)-(60b) is well-posed, and the solution $(\boldsymbol{\sigma}_{h},\boldsymbol{u}_{h},\phi_{h})\in{\Sigma}^{\rm tn}_{h}\times\mathring{\mathbb{V}}^{{\rm curl\,}}_{(k,\ell),h}\times\mathring{\mathbb{V}}^{{\rm grad\,}}_{\ell+1,h}$ coincides with the mixed finite element method (39a)-(39b).
Proof.
For $\boldsymbol{\tau}_{h}\in\Sigma_{h}^{-1}$ and $\psi_{h}\in\mathring{\mathbb{V}}_{h}^{{\rm grad\,}}$ satisfying
$b_{h}(\boldsymbol{\tau}_{h},\psi_{h};\boldsymbol{v}_{h})+c_{h}(\boldsymbol{\tau}_{h},\boldsymbol{\mu}_{h})=0,\quad\boldsymbol{v}_{h}\in\mathring{\mathbb{V}}_{h}^{{\rm curl\,}},\boldsymbol{\mu}_{h}\in\Lambda_{h},$
we have $\boldsymbol{\tau}_{h}\in{\Sigma}^{\rm tn}_{h}$, and $b_{h}(\boldsymbol{\tau}_{h},\psi_{h};\boldsymbol{v}_{h})=0$ for $\boldsymbol{v}_{h}\in\mathring{\mathbb{V}}_{h}^{{\rm curl\,}}.$ By (44), we obtain the discrete coercivity $\|\boldsymbol{\tau}_{h}\|^{2}+|\psi_{h}|_{1}^{2}\lesssim\|\boldsymbol{\tau}_{h}\|^{2}.$ Then we get the well-posedness of the hybridized mixed finite element method (60a)-(60b) by applying the Babuška-Brezzi theory [5] with the discrete inf-sup condition (LABEL:hybridizationinfsup).
By (60b) with $\boldsymbol{v}_{h}=0$, $\boldsymbol{\sigma}_{h}\in{\Sigma}^{\rm tn}_{h}$, thus $(\boldsymbol{\sigma}_{h},\boldsymbol{u}_{h},\phi_{h})$ satisfies (39a)-(39b). ∎
Theorem 5.3.
Let $(\boldsymbol{\sigma},0,\boldsymbol{u})$ and $(\boldsymbol{\sigma}_{h},0,\boldsymbol{u}_{h},\boldsymbol{\lambda}_{h})$ be the solution of the mixed formulation (34a)-(34b) and the mixed finite element method (60) respectively. Assume $\boldsymbol{\sigma}\in H^{k}(\Omega;\mathbb{T})$ and $\boldsymbol{u},{\rm curl\,}\boldsymbol{u}\in H^{k}(\Omega;\mathbb{R}^{3})$. Then
(63) $\left\|\boldsymbol{n}_{F}\times{\rm curl\,}\boldsymbol{u}_{h}-\boldsymbol{\lambda}_{h}\right\|_{-1/2,h}\lesssim h^{k}(h|\boldsymbol{\sigma}|_{k}+h|\boldsymbol{u}|_{k}+|{\rm curl\,}\boldsymbol{u}|_{k}).$
Proof.
By the proof of the discrete inf-sup condition (LABEL:hybridizationinfsup),
$\left\|\boldsymbol{n}_{F}\times{\rm curl\,}(I_{h}^{{\rm curl\,}}\boldsymbol{u}-\boldsymbol{u}_{h})-(Q_{\mathcal{F}_{h}}^{k-1}(\boldsymbol{n}_{F}\times{\rm curl\,}\boldsymbol{u})-\boldsymbol{\lambda}_{h})\right\|_{1/2,h}$
$\lesssim$ $\sup_{\boldsymbol{\tau}_{h}\in\Sigma_{k-1,h}^{-1}}\frac{b_{h}(\boldsymbol{\tau}_{h},0;I_{h}^{{\rm curl\,}}\boldsymbol{u}-\boldsymbol{u}_{h})+c_{h}(\boldsymbol{\tau}_{h},Q_{\mathcal{F}_{h}}^{k-1}(\boldsymbol{n}_{F}\times{\rm curl\,}\boldsymbol{u})-\boldsymbol{\lambda}_{h})}{\left\|\boldsymbol{\tau}_{h}\right\|}.$
Thanks to (24) and (60a), we have
$b_{h}(\boldsymbol{\tau}_{h},0;I_{h}^{{\rm curl\,}}\boldsymbol{u}-\boldsymbol{u}_{h})+c_{h}(\boldsymbol{\tau}_{h},Q_{\mathcal{F}_{h}}^{k-1}(\boldsymbol{n}_{F}\times{\rm curl\,}\boldsymbol{u})-\boldsymbol{\lambda}_{h})=(\boldsymbol{\sigma}_{h}-\boldsymbol{\sigma},\boldsymbol{\tau}_{h}).$
Hence
$\left\|\boldsymbol{n}_{F}\times{\rm curl\,}(I_{h}^{{\rm curl\,}}\boldsymbol{u}-\boldsymbol{u}_{h})-(Q_{\mathcal{F}_{h}}^{k-1}(\boldsymbol{n}_{F}\times{\rm curl\,}\boldsymbol{u})-\boldsymbol{\lambda}_{h})\right\|_{1/2,h}\lesssim\|\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h}\|.$
Therefore we can derive estimate (63) from (48) and the error estimate of $I_{h}^{{\rm curl\,}}$. ∎
5.2. Connection to other methods
The pair $(\boldsymbol{v}_{h},\boldsymbol{\mu}_{h})$ can be understood as a weak function. Define $({\rm grad\,}{\rm curl\,})_{w}:\mathring{\mathbb{V}}_{h}^{{\rm curl\,}}\times\Lambda_{h}\rightarrow{\Sigma}_{h}^{-1}$ by
$\left(({\rm grad\,}{\rm curl\,})_{w}(\boldsymbol{v}_{h},\boldsymbol{\mu}_{h}),\boldsymbol{\tau}_{h}\right)=-b_{h}\left(\boldsymbol{\tau}_{h},0;\boldsymbol{v}_{h}\right)-c_{h}(\boldsymbol{\tau}_{h},\boldsymbol{\mu}_{h}),\quad\boldsymbol{\tau}_{h}\in{\Sigma}_{k-1,h}^{-1}.$
By (60a), we can eliminate $\boldsymbol{\sigma}_{h}$ element-wise and write $\boldsymbol{\sigma}_{h}=({\rm grad\,}{\rm curl\,})_{w}(\boldsymbol{u}_{h},\boldsymbol{\lambda}_{h})$. Then the hybridized mixed finite element method (60a)-(60b) can be recast as: find $(\boldsymbol{u}_{h},\lambda_{h},\phi_{h})\in\mathring{\mathbb{V}}_{(k,\ell),h}^{{\rm curl\,}}\times\Lambda_{k-1,h}\times\mathring{\mathbb{V}}_{\ell+1,h}^{{\rm grad\,}}$ such that
$(({\rm grad\,}{\rm curl\,})_{w}(\boldsymbol{u}_{h},\boldsymbol{\lambda}_{h}),({\rm grad\,}{\rm curl\,})_{w}(\boldsymbol{v}_{h},\boldsymbol{\mu}_{h}))-(\boldsymbol{v}_{h},{\rm grad\,}\phi_{h})$ $=\langle\boldsymbol{f},\boldsymbol{v}_{h}\rangle,$
$(\boldsymbol{u}_{h},{\rm grad\,}\psi_{h})$ $=0,$
for all $(\boldsymbol{v}_{h},\mu_{h},\psi_{h})\in\mathring{\mathbb{V}}_{(k,\ell),h}^{{\rm curl\,}}\times\Lambda_{k-1,h}\times\mathring{\mathbb{V}}_{\ell+1,h}^{{\rm grad\,}}$. That is we obtain a stabilization free weak Galerkin method for solving the quad-curl problem. The stabilization free is due to the inf-sup condition (62). Indeed, the inf-sup condition (62) is equivalent to
$\|{\rm grad\,}_{\mathcal{T}_{h}}{\rm curl\,}\boldsymbol{v}_{h}\|+\left\|\boldsymbol{n}_{F}\times{\rm curl\,}\boldsymbol{v}_{h}-\boldsymbol{\mu}_{h}\right\|_{1/2,h}\lesssim\|({\rm grad\,}{\rm curl\,})_{w}(\boldsymbol{v}_{h},\boldsymbol{\mu}_{h})\|$
for $\boldsymbol{v}_{h}\in\mathring{\mathbb{V}}_{(k,\ell),h}^{{\rm curl\,}}$ and $\boldsymbol{\mu}_{h}\in\Lambda_{k-1,h}$. This means $\|({\rm grad\,}{\rm curl\,})_{w}(\boldsymbol{v}_{h},\boldsymbol{\mu}_{h})\|$ is a norm on space $K_{h}^{c}\times\Lambda_{k-1,h}$.
For the hybridized mixed finite element method (60a)-(60b) of the lowest order $k=1$ and $\ell=0,1$, it is also related to a nonconforming finite element method.
We first recall the $H({\rm grad\,}{\rm curl\,})$ nonconforming finite elements constructed in [28, 45]. The space of shape functions is ${\rm grad\,}\mathbb{P}_{\ell+1}(T)\oplus(\boldsymbol{x}-\boldsymbol{x}_{T})\times\mathbb{P}_{1}\left(T;\mathbb{R}^{3}\right)$ for $\ell=0,1$. The degrees of freedom are given by
(64a) $\int_{e}\boldsymbol{v}\cdot\boldsymbol{t}\,q\,{\rm d}s,$ $\;\quad q\in\mathbb{P}_{\ell}(e)\text{ on each }e\in\mathcal{E}(T),$
[…]
Define the global $H({\rm grad\,}{\rm curl\,})$-nonconforming element space
$W_{h}:=$$\{\boldsymbol{v}_{h}\in{L}^{2}(\Omega;\mathbb{R}^{3}):\,\boldsymbol{v}_{h}|_{T}\in{\rm grad\,}\mathbb{P}_{\ell+1}(T)\oplus(\boldsymbol{x}-\boldsymbol{x}_{T})\times\mathbb{P}_{1}\left(T;\mathbb{R}^{3}\right),\ \forall~{}T\in\mathcal{T}_{h},$
[…]
The interpolation operator $I_{h}^{{\rm curl\,}}$ can be extended to $W_{h}$, which is well-defined. We have $I_{h}^{{\rm curl\,}}W_{h}=\mathring{\mathbb{V}}_{h}^{{\rm curl\,}}$, and (24) still holds for $\boldsymbol{v}\in W_{h}$.
An $H({\rm grad\,}{\rm curl\,})$-nonconforming finite element method for quad-curl problem (32) is to find $\boldsymbol{w}_{h}\in W_{h}$ and $\phi_{h}\in\mathring{\mathbb{V}}_{h}^{{\rm grad\,}}$ such that
(65a) $({\rm grad\,}_{\mathcal{T}_{h}}{\rm curl\,}_{\mathcal{T}_{h}}\boldsymbol{w}_{h},{\rm grad\,}_{\mathcal{T}_{h}}{\rm curl\,}_{\mathcal{T}_{h}}\boldsymbol{v}_{h})-(I_{h}^{{\rm curl\,}}\boldsymbol{v}_{h},{\rm grad\,}\phi_{h})$ $=\langle\boldsymbol{f},I_{h}^{{\rm curl\,}}\boldsymbol{v}_{h}\rangle,$
(65b) $(I_{h}^{{\rm curl\,}}\boldsymbol{w}_{h},{\rm grad\,}\psi_{h})$ $=0,$
for all $\boldsymbol{v}_{h}\in W_{h}$ and $\psi_{h}\in\mathring{\mathbb{V}}_{h}^{{\rm grad\,}}$. Nonconforming finite element method (65a)-(65b) is a modification of those in [28, 45] by introducing interpolation operator $I_{h}^{{\rm curl\,}}$. In other words, we identify the complex that accommodates the non-conforming finite elements constructed in [28, 45] and generalize these elements to arbitrary orders.
Lemma 5.4.
For $\boldsymbol{v}_{h}\in W_{h}$ satisfying $(I_{h}^{{\rm curl\,}}\boldsymbol{v}_{h},{\rm grad\,}q_{h})=0$ for all $q_{h}\in\mathring{\mathbb{V}}_{h}^{{\rm grad\,}}$, we have the discrete Poincaré inequality
(66) $\|\boldsymbol{v}_{h}\|+\|{\rm curl\,}_{\mathcal{T}_{h}}\boldsymbol{v}_{h}\|\lesssim\|{\rm grad\,}_{\mathcal{T}_{h}}{\rm curl\,}_{\mathcal{T}_{h}}\boldsymbol{v}_{h}\|.$
Proof.
By (4.12) in [28], it follows
$\|\boldsymbol{v}_{h}-I_{h}^{{\rm curl\,}}\boldsymbol{v}_{h}\|\lesssim h\|{\rm curl\,}_{\mathcal{T}_{h}}\boldsymbol{v}_{h}\|.$
Thanks to the discrete Poincaré inequality for space $\mathring{\mathbb{V}}_{h}^{{\rm curl\,}}$, we have
$\|I_{h}^{{\rm curl\,}}\boldsymbol{v}_{h}\|\lesssim\|{\rm curl\,}(I_{h}^{{\rm curl\,}}\boldsymbol{v}_{h})\|.$
Combining the last two inequalities gives
$\|\boldsymbol{v}_{h}\|\leq\|\boldsymbol{v}_{h}-I_{h}^{{\rm curl\,}}\boldsymbol{v}_{h}\|+\|I_{h}^{{\rm curl\,}}\boldsymbol{v}_{h}\|\lesssim h\|{\rm curl\,}_{\mathcal{T}_{h}}\boldsymbol{v}_{h}\|+\|{\rm curl\,}(I_{h}^{{\rm curl\,}}\boldsymbol{v}_{h})\|.$
By the commutative property of $I_{h}^{{\rm curl\,}}$, it holds that
$\|{\rm curl\,}(I_{h}^{{\rm curl\,}}\boldsymbol{v}_{h})\|\lesssim\|{\rm grad\,}_{\mathcal{T}_{h}}{\rm curl\,}_{\mathcal{T}_{h}}\boldsymbol{v}_{h}\|.$
Hence
$\|\boldsymbol{v}_{h}\|\lesssim\|{\rm curl\,}_{\mathcal{T}_{h}}\boldsymbol{v}_{h}\|+\|{\rm grad\,}_{\mathcal{T}_{h}}{\rm curl\,}_{\mathcal{T}_{h}}\boldsymbol{v}_{h}\|.$
Finally apply the discrete Poincaré inequality for $H^{1}$-nonconforming linear element to end the proof. ∎
Using the discrete Poincaré inequality (66), we have the well-posedness.
Lemma 5.5.
Nonconforming finite element method (65a)-(65b) is well-posed.
We will show the equivalence between the hybridized mixed finite element method (60a)-(60b) with $k=1$ and nonconforming finite element method (65a)-(65b).
Theorem 5.6.
Let $(\boldsymbol{w}_{h},\phi_{h})\in W_{h}\times\mathring{\mathbb{V}}_{h}^{{\rm grad\,}}$ be the solution of the nonconforming finite element method (65a)-(65b). Then $({\rm grad\,}_{\mathcal{T}_{h}}{\rm curl\,}_{\mathcal{T}_{h}}\boldsymbol{w}_{h},I_{h}^{{\rm curl\,}}\boldsymbol{w}_{h},\phi_{h},Q_{\mathcal{F}_{h}}(\boldsymbol{n}_{F}\times{\rm curl\,}\boldsymbol{w}_{h}))\in{\Sigma}_{0,h}^{-1}\times\mathring{\mathbb{V}}^{{\rm curl\,}}_{h}\times\mathring{\mathbb{V}}^{{\rm grad\,}}_{h}\times\Lambda_{h}$ is the solution of the hybridized mixed finite element method (60a)-(60b) with $k=1$.
Proof.
Choose $\boldsymbol{v}_{h}\in W_{h}$ such that DoF (64a) vanishes, then $I_{h}^{{\rm curl\,}}\boldsymbol{v}_{h}=0$. Applying the integration by parts on the left hand side of (65a), we get
00 0 $=\sum_{T\in\mathcal{T}_{h}}(({\rm grad\,}{\rm curl\,}\boldsymbol{w}_{h})\boldsymbol{n},{\rm curl\,}\boldsymbol{v}_{h})_{\partial T}$
$=\sum_{T\in\mathcal{T}_{h}}(\boldsymbol{n}\times({\rm grad\,}{\rm curl\,}\boldsymbol{w}_{h})\boldsymbol{n},\boldsymbol{n}\times{\rm curl\,}\boldsymbol{v}_{h})_{\partial T}$
$=\sum_{F\in\mathring{\mathcal{F}}_{h}}(\llbracket\boldsymbol{n}\times({\rm grad\,}_{\mathcal{T}_{h}}{\rm curl\,}_{\mathcal{T}_{h}}\boldsymbol{w}_{h})\boldsymbol{n}\rrbracket,\boldsymbol{n}_{F}\times{\rm curl\,}_{\mathcal{T}_{h}}\boldsymbol{v}_{h})_{F}.$
By the arbitrariness of the DoF (64b) for $\boldsymbol{v}_{h}$, we obtain $\llbracket\boldsymbol{n}\times({\rm grad\,}_{\mathcal{T}_{h}}{\rm curl\,}_{\mathcal{T}_{h}}\boldsymbol{w}_{h})\boldsymbol{n}\rrbracket_{F}=0$ for all $F\in\mathring{\mathcal{F}}_{h}$, that is ${\rm grad\,}_{\mathcal{T}_{h}}{\rm curl\,}_{\mathcal{T}_{h}}\boldsymbol{w}_{h}\in\Sigma_{h}^{\rm tn}$. For all $\boldsymbol{v}_{h}\in W_{h}$ and $\boldsymbol{\mu}_{h}\in\Lambda_{h}$, we get from (24), the integration by parts, and (65a) that
$\quad b_{h}({\rm grad\,}_{\mathcal{T}_{h}}{\rm curl\,}_{\mathcal{T}_{h}}\boldsymbol{w}_{h},\phi_{h};I_{h}^{{\rm curl\,}}\boldsymbol{v}_{h})+c_{h}({\rm grad\,}_{\mathcal{T}_{h}}{\rm curl\,}_{\mathcal{T}_{h}}\boldsymbol{w}_{h},\boldsymbol{\mu}_{h})$
$=b_{h}({\rm grad\,}_{\mathcal{T}_{h}}{\rm curl\,}_{\mathcal{T}_{h}}\boldsymbol{w}_{h},\phi_{h};I_{h}^{{\rm curl\,}}\boldsymbol{v}_{h})$
$=b_{h}({\rm grad\,}_{\mathcal{T}_{h}}{\rm curl\,}_{\mathcal{T}_{h}}\boldsymbol{w}_{h},0;\boldsymbol{v}_{h})+(I_{h}^{{\rm curl\,}}\boldsymbol{v}_{h},{\rm grad\,}\phi_{h})$
$=-({\rm grad\,}_{\mathcal{T}_{h}}{\rm curl\,}_{\mathcal{T}_{h}}\boldsymbol{w}_{h},{\rm grad\,}_{\mathcal{T}_{h}}{\rm curl\,}_{\mathcal{T}_{h}}\boldsymbol{v}_{h})+(I_{h}^{{\rm curl\,}}\boldsymbol{v}_{h},{\rm grad\,}\phi_{h})=-\langle\boldsymbol{f},I_{h}^{{\rm curl\,}}\boldsymbol{v}_{h}\rangle.$
Notice that $I_{h}^{{\rm curl\,}}:W_{h}\to\mathring{\mathbb{V}}_{h}^{{\rm curl\,}}$ is onto, hence $({\rm grad\,}_{\mathcal{T}_{h}}{\rm curl\,}_{\mathcal{T}_{h}}\boldsymbol{w}_{h},\phi_{h})$ satisfies (60b).
On the side hand, for all $\boldsymbol{\tau}_{h}\in\Sigma_{0,h}^{-1}$ and $\psi_{h}\in\mathring{\mathbb{V}}_{h}^{{\rm grad\,}}$, apply (24), (65b) and the integration by parts to get
$\quad({\rm grad\,}_{\mathcal{T}_{h}}{\rm curl\,}_{\mathcal{T}_{h}}\boldsymbol{w}_{h},\boldsymbol{\tau}_{h})+b_{h}\left(\boldsymbol{\tau}_{h},\psi_{h};I_{h}^{{\rm curl\,}}\boldsymbol{w}_{h}\right)+c_{h}(\boldsymbol{\tau}_{h},Q_{\mathcal{F}_{h}}(\boldsymbol{n}_{F}\times{\rm curl\,}\boldsymbol{w}_{h}))$
$=({\rm grad\,}_{\mathcal{T}_{h}}{\rm curl\,}_{\mathcal{T}_{h}}\boldsymbol{w}_{h},\boldsymbol{\tau}_{h})+b_{h}\left(\boldsymbol{\tau}_{h},0;\boldsymbol{w}_{h}\right)+c_{h}(\boldsymbol{\tau}_{h},\boldsymbol{n}_{F}\times{\rm curl\,}\boldsymbol{w}_{h})$
$\quad+(I_{h}^{{\rm curl\,}}\boldsymbol{w}_{h},{\rm grad\,}\psi_{h})$
$=({\rm grad\,}_{\mathcal{T}_{h}}{\rm curl\,}_{\mathcal{T}_{h}}\boldsymbol{w}_{h},\boldsymbol{\tau}_{h})+b_{h}\left(\boldsymbol{\tau}_{h},0;\boldsymbol{w}_{h}\right)-\sum_{T\in\mathcal{T}_{h}}\left(\boldsymbol{n}\times\boldsymbol{\tau}_{h}\boldsymbol{n},\boldsymbol{n}\times{\rm curl\,}\boldsymbol{w}_{h}\right)_{\partial T}=0.$
That is $({\rm grad\,}_{\mathcal{T}_{h}}{\rm curl\,}_{\mathcal{T}_{h}}\boldsymbol{w}_{h},I_{h}^{{\rm curl\,}}\boldsymbol{w}_{h},Q_{\mathcal{F}_{h}}(\boldsymbol{n}_{F}\times{\rm curl\,}\boldsymbol{w}_{h}))$ satisfies (60a). ∎